### v1

# How Books Tell a History of Statistics in Portugal: Works of Foreigners, $\boldsymbol{Estrangeirados}$, and Others

How Books Tell a History of Statistics in Portugal:
Works of Foreigners, $\boldsymbol{Estrangeirados}$, and Others
Dinis Pestana1,3,4, Rui Santos2,3
1 DEIO, Faculdade de Ciências da Universidade de Lisboa (FCUL)
2 Escola Superior de Tecnologia e Gestão, Instituto Politécnico de Leiria
3 Centro de Estatística e Aplicações (CEA/UL)
4 Instituto de Investigação Científica Bento da Rocha Cabral
  • [Uncaptioned image]
  • Dedicated to Fernando Rosado, in gratitude for his outstanding edition of the Boletim da Sociedade Portuguesa de Estatística and the Memorial da Sociedade Portuguesa de Estatística.
  • [Uncaptioned image]
“Statistics is the grammar of science.”
Karl Pearson
1 Introduction
Sketching the History of Statistics in Portugal naturally requires discussing the influence of foreigners and “$estrangeirados$”, foreign-influenced individuals(1)(1)(1)  In the 17th and 18th centuries, some Portuguese intellectuals, who were exposed to scientific innovations and the Enlightenment philosophy abroad, had a far-reaching influence on the modernization of culture in Portugal. Among these “$Estrangeirados$” (as they are referred to in Portuguese history) are Ribeiro Sanches, Luís António Verney, Francisco Xavier de Oliveira (Cavaleiro d’Oliveira), D. Luís da Cunha (whose political testament recommended Sebastião José de Carvalho e Melo, later Marquês de Pombal, to King D. José), and the Marquês de Pombal himself. The Pombaline reform of the university in 1772 was certainly due to the influential actions of the $estrangeirados$.
The liberal struggles of the 19th century forced many illustrious Portuguese to exile, particularly in the United Kingdom and/or France. Some of these new estrangeirados, namely Solano Constâncio, Alexandre Herculano, and Oliveira Marreca, were important agents for the dissemination of Statistics in Portugal. The politician Passos Manuel, exiled in La Coruña-Plymouth-Belgium-Paris between 1828 and 1832, carried out important reforms during his brief tenure as Minister of the Kingdom in 1836-1837, both in secondary and higher education. One of the consequences of his reform was the creation of a course of Economia Política [Political Economy], whose Professor, Adrião Forjaz de Sampaio, included the teaching of Princípios de Statistica [Statistical Primer] in its program from 1841 onwards. We use the term $estrangeirado$ in a very general sense: estrangeirados will also be much later Zaluar Nunes (first as a student in Paris, then as a political exile in Brazil, due to the diaspora of intellectuals persecuted by Salazarism), Bento Murteira in London, Gustavo de Castro in Paris, Sebastião e Silva in Italy, Dias Agudo, Tiago de Oliveira, and Pedro Braumann in the USA, and later many scholarship holders who pursued doctorates abroad.
, without neglecting the role of “others”. Plagiating Mae West, who when asked what kind of men she preferred, candidly replied: “Personally, I like two types of men — domestic and foreign”, in our opinion many “Others”, who hadn’t the opportunity of extended partnerships abroad, had also a relevant role in the development of Mathematical Statistics in Portugal, and their contributions must be recognised and praised as much as those of the estrangeirados. Notorious examples are António Simões Neto, a man of culture and one of the early enthusiasts of Bayesian ideas in Portugal, who influenced many students at the Faculty of Sciences of Lisbon (namely Dinis Pestana and Maria Antónia Amaral Turkman), Fernando Rosado (to whom we dedicate this sketch of the history of Statistics in Portugal, thanking his contribution to a “Portugaliae Monumenta Statistica” with his fine edition of the Boletim da Sociedade Portuguesa de Estatística [Bulletin of the Portuguese Statistical Society]), Salomé Cabral or Manuela Neves (recently recipient of the Portuguese Statistical Society Life Achievement Award), whose doctoral dissertations were supervised by Tiago de Oliveira, or Rita Vasconcelos, who contributed so much for the progress of the Probability and Statistics group at the University of Madeira.
Regarding estrangeirados, there is quite a bit of information, but when it comes to the influence of foreigners, the gaps are significant, especially considering the enormous long-distance influence of works already mentioned in the 19th century (Quételet by Forjaz de Sampaio, Bertrand and Poincaré in Sidónio Paes’s thesis) or used as recommended bibliography in courses or research seminars (for example, books by Fisher, Kendall and Stuart, Cramér, Loève, Marek Fisz, Feller, Gnedenko, Cox, Zar, Rohatgi, …).
Information is also quite lacking about the actual role of foreigners who came to Portugal as invited congress participants, and even more so about those who came on their own initiative as participants. There is also little information on the relevance to the development of Statistics in Portugal of those who served on advisory boards of research centers.
On the other hand, it is interesting to list which foreign statisticians were awarded honorary doctorates by Portuguese universities, which foreigners chose to spend part of their academic life in Portugal, and which books were translated and served as useful aids in the teaching, research, and dissemination of Statistics in Portugal.
Furthermore, research in university libraries and the National Library allowed us to identify and consult a reasonable number of statistical works that have been published, providing a reliable indicator of how the introduction of Mathematical Statistics concepts and the evolution of their teaching and dissemination progressed.
Although we are aware that we are far from an exhaustive presentation, we will attempt to give a coordinated overview of how the teaching and dissemination of Statistics developed in the 19th century (Section 2). During this period, the increasing importance of official statistics and the introduction of Probability teaching at the Polytechnic School contributed to the recognition of Statistics as a mathematical discipline, culminating with Sidónio Paes submission of the first doctoral thesis in the field of Mathematical Statistics in Portugal, at the Faculty of Mathematics of the University of Coimbra.
Concerning the advancements in Mathematical Statistics in the 20th century up to 1974 (Section 3), we highlight how the teaching of Probability, particularly at the Faculdade de Ciências [Faculty of Sciences], produced excellent professors such as Varennes e Mendonça, Zaluar Nunes, or Sebastião e Silva for Instituto Superior de Agronomia (ISA) [Higher Institute of Agronomy], and Leite Pinto in ISEG. These professors, and Bento Murteira, who worked under Maurice Kendall in London, imported the new statistical concepts, particularly those of Pearson (father and son) and Fisher. Fisher’s influence was significant at ISA, but we should not overlook the potential role the Institute of Anthropology in Coimbra could have played due to the hiring of Stevens, a disciple of Fisher, who published a series of remarkable “Methodological Issues” in the Revista da Faculdade de Ciências da Universidade de Coimbra [Review of the Faculty of Sciences of the University of Coimbra]. Finally, we include some informal notes about the last fifty years (Section 4), for which it seems premature to adopt a historical perspective.
2 19th Century
The Venetian geographer and statistician Adriano Balbi (1782-1848) was in Portugal in 1820, before settling in Paris from 1821 to 1832, where he changed his name to Adrien Balbi. It was in Paris, in 1822, that he published the two volumes of the Essai statistique sur le royaume de Portugal et d’Algarve [Statistical Essay on the Kingdom of Portugal and Algarve] and also Variétés politico-statistiques de la monarchie portugaise [Various Political-Statistical Observations on the Portuguese Monarchy].
Refer to caption Refer to caption Refer to caption
(a) Essai Statistique (b) Variétés Politico-Statistiques (c) Critical remarks
Figure 1: Covers of Balbi’s books on the Kingdom of Portugal and Villela da Silva’s Critiques
Some statements in Balbi’s books probably irritated the portuguese establishment, since in 1828 Villela da Silva published negative critical observations on Balbi’s works at the Portuguese Royal Press.
On the other hand, Balbi’s work was highly praised by Solano Constâncio (1777-1846), who in 1822 published an analysis of Variétés politico-statistiques sur la monarchie portugaise in the Anais das Ciências, Artes, e Letras [Annals of Sciences, Arts, and Letters], Volume XVII, 84–103.(2)(2)(2)  Pages 212–222 of the 1995 reedition of Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], a selection of works by Solano Constâncio [1], published by the Bank of Portugal, edited by José Luís Cardoso, accessible at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf. The pages cited ahead are from this edition. It should be noted, incidentally, that Solano Constâncio, editor of the above mentioned Anais das Ciências, Artes, e Letras, in this critical review refers to the article Considerações sobre a Estatística [Considerations on Statistics] by Cândido Xavier, published in 1820, in Volume X, pp. 134–172.
Solano Constâncio was a remarkable expatriate polygraph, who studied medicine in the United Kingdom between 1791 and 1797, later living in Paris from 1797 to 1799. He returned to Portugal, but in 1807 emigrated again to Paris. From 1822 to 1826 he lived in the United States of America, then returned to Paris, where he died in 1846. Machado de Sousa (1979) [2] and Cardoso (1995) [3] are interesting sources of information about the wide range of interests of this remarkable intellectual.
Regarding Statistics, in the “Preliminary Discourse” of the Anais das Ciências, das Artes e das Letras, Volume I (1818), pp. 1-37, (pp. 81-99 of Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Solano Constâncio states that
Statistics, a science equally of our times, will also provide material for some articles and for examining the most notable works or news appearing on this subject.” (p. 94)
and on p. 176 of the Anais das Ciências, das Artes e das Letras, in the critical review of An inquiry concerning population, etc…, ou investigação acerca da população e da faculdade de multiplicação da espécie humana; obra destinada a refutar a doutrina do ensaio de M. Malthus sobre esse assunto por W. Godwin, Londres, 1820 [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr. Malthus’s essay on this subject by W. Godwin, London, 1820] (Volume XII (1821), 63–104.):
All the statistical notions that have been collected for a century agree in showing the great uniformity of nature’s laws regarding reproduction of the species. The proportion of births and deaths, the number of males and females, the proportion of children per marriage, the number of marriages relative to the total population, and the age proportions, in historically regulated states, are facts that, the more they have been observed, the more conformity has been found between regions, climates, and diverse peoples. Despite the great imperfection of statistical works throughout Europe, those countries where censuses and population maps, birth and death registers have been pursued longer and with more care yield very analogous results among themselves. Among many other important inferences, it uniformly results from all these documents that the human species has a very small tendency to increase regularly and permanently.
The Armazém de Conhecimentos Úteis, nas Artes e Ofícios, ou Colecção de Tratados, Receitas e Invenções de Utilidade Geral [Repository of Useful Knowledge, in Arts and Crafts, or Collection of Treatises, Recipes, and Inventions of General Utility], J.-P. Aillaud, Paris, 1838, contains “Considerações sobre a estatística” [“Considerations on Statistics”] (1, 51–64), reproduced on pp. 271–277 of Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy]; we quote from p. 276:
induction from statistical maps will prove the notable utility of this science, which is now indispensable for governing and enlightening the public.
We find it interesting to note that in the 1821 citation the refutation of Malthus’ exponential population growth, concluding that the human species has very little tendency to increase regularly and permanently foreshadows what Verhulst [4, 5, 6] would later formalize, in 1837, with the logistic model; and that, in the 1838 citation, the notable utility of this science, which is now indispensable for […] enlightening the public has a meaning very close to the famous Thomas Carlyle’s [7] statement in Chap. II of Chartism: “A judicious man uses statistics, not to get knowledge, but to save himself from having ignorance foisted upon him.”.
More interesting still is the use in those excerpts of the words “inferences” and “induction”. For Solano Constâncio, Statistics has an inferential component that goes beyond its merely descriptive component. It is perhaps for this reason that he distinguishes “Estatística”  [Statistics] from “Estadística” [Stadistics], etymogically derived from “Estado” (government), the science of governing or dealing with political affairs, more connected to what is termed official statistics.
It was Francisco Solano Constâncio who, in 1836, in his Novo Diccionario Critico e Etymologico da Lingua Portugueza [New Critical and Etymological Dictionary of the Portuguese Language], defined, for the first time in Portuguese, Statistics as the science “that deals with the enumeration of everything that forms the strength of a nation”. He added a list of subjects: population, agriculture, industry, education, public revenues and expenses, military strength, and property distribution. The Novo Diccionario Critico e Etymologico da Lingua Portugueza achieved enduring editorial success, with 11 editions until 1877.
Refer to caption Refer to caption Refer to caption
(a) Diccionario’s cover (b) entry “ESTATISTICA” (c) page 66 of the Diccionario
Figure 2: Solano Constâncio’ s Novo Diccionario Critico e Etymologico da Lingua Portugueza
One of the consequences of the 1836 reform by Passos Manuel was the creation of a Chair of Political Economy at the Faculty of Law (resulting from merging the faculties of Laws and of Canons) of the University of Coimbra. Adrião Forjaz de Sampaio included in the program of the chair an informal presentation of Descriptive Statistics, from the perspective of preparing official statistics, heavily influenced by the journal of works of the Société Française de Statistique Universelle (Fontes de Sousa, 2005 [8]; Ferrão, 2006 [9]).
  • [Uncaptioned image]
  • Thus, the pioneer of Statistics teaching at the Faculty of Law of the University of Coimbra was Adrião Forjaz de Sampaio (1810-1874). He published Primeiros Elementos da Sciencia da Statistica [First Elements of the Science of Statistics], in which the influence of Quételet is noted. This manual presents the elements of Descriptive Statistics, giving some prominence to averages and dispersions, with rudimentary mathematical formalism. It is interesting to note that in the title, in later editions Forjaz de Sampaio uses “Stadistics” instead of “Statistics”, and that in 1857 José Dias Ferreira (1837-1907), then still a student, published at the University Press an Ensaio sobre os Primeiros Elementos da Theoria da Estadística do Exmo. Senhor Adrião Pereira Forjaz de Sampaio [Essay on the First Elements of the Theory of Statistics by the Honorable Adrião Pereira Forjaz de Sampaio].
In 1887 José Frederico Laranjo published a new program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra, “with the approval of Dr. Manuel Nunes Giraldes”. Citing Bastien (1997), [10] “Note that the Chair of Political Economy, after its creation in 1836, and until 1871, was mostly held by Adrião Forjaz de Sampaio and after that date, and until 1899, by Manuel Nunes Giraldes. Despite the strong relationship of Frederico Laranjo with the Chair of Political Economy, particularly in terms of teaching, program development, publication of teaching manuals, and doctoral thesis examinations […] he was never awarded the respective chair.”. The lessons of J. F. Laranjo in Princípios de Economia Política [Principles of Political Economy] (1891), reprinted in [11] do not contain notes on Statistics.
The dichotomy between Estatística/Estadística [Statistics/Stadistics] continued to divide those interested in this new science(s) in the 19th century. On March 26, 1852, Alexandre Herculano(3)(3)(3)  A remarkable historian, author of an excellent four volumes very comprehensive História de Portugal [History of Portugal] and of História da Origem e Estabelecimento da Inquisição em Portugal. [History of the Origin and Establishment of the Inquisition in Portugal]. He compiled and edited Portugaliae Monumenta Historica, and was also a highly praised romantic novelist author namely of historical novels. (1810-1877) — who was exiled in England and then in France between 1831 and 1832 — proposed to the Royal Academy of Sciences that the section of Economic and Administrative Sciences should “draft the necessary instructions and a series of statistical questions, in harmony with the current state of science, which the Academy should bring to the Government’s attention, seeking to obtain from it an order for civil, ecclesiastical, and military officials, as well as judicial magistrates and elected authorities of any order and denomination, to respond within the scope of their respective actions to the aforementioned questions, with these responses being transmitted to the Academy.”
This interest of Herculano in Statistics was not an isolated episode in his vast array of interests and activities. Two excerpts from Emigration, in Opúsculos, Volume IV [12] show that Herculano maintained his interest in Statistics, appreciating its contribution to the objective understanding of social phenomena:
We thus have a set of respectable opinions on the insufficiency of rural wages. These opinions are fully confirmed by statistics, revealing with the irresistible eloquence of numbers the true situation of the laborer, both in terms of his resources and his needs. Incomplete for covering only part of the districts of the kingdom, deficient due to omissions and lack of specification in the elements provided by some municipalities, the statistical tables added to the Parliamentary Inquiry are still numerous enough in their various types to allow general conclusions to be drawn from them.
If the rejection of official means of information seems inconvenient and unfair to me, I find the absolute condemnation of statistics itself, voted to ostracism as a conjectural science, even more peculiar. There is necessarily a misunderstanding here. Statistics aims to collect and methodically organize social facts that can be expressed in numbers. Nothing less conjectural. If, instead of facts, assumptions are collected, that is not statistics, it is fiction. In applied statistics, conjectures and the errors often derived from them are possible; but to infer from this the futility of statistics is the same as denying the scientific validity of arithmetic because sums or multiplications are often mistaken. Generally, what do the laws and measures of a generic nature, whether of a legal, moral, or economic order, depend on? They depend on a correlative statistical study. It is in this study that their raison d’être lies.
António de Oliveira Marreca (1805-1889) also went into exile in England and France, returning to Portugal in 1833. He was highly influential in the institutionalization of Political Economy in Portugal, for detailed information cf.  Veríssimo Serrão (1978) [13] and Almodôvar (1995) [14].
  • [Uncaptioned image]
  • In Parecer e Memoria sobre a proposta, que apresentou o Sr. Alexandre Herculano: para que a secção de sciencias económicas e administrativas redigisse um Projecto de Estadística [A Review and Memoir on Mr. Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Statistics], Typographia da Academia Real das Sciencias, 1854, Oliveira Marreca returns to the use of the terminology $estadistica$.
Oliveira Marreca was one of the administrators of the Imprensa Nacional de Lisboa [National Press of Lisbon] from 1834 to 1838. Arguably Oliveira Marreca influenced the publication at the Imprensa Nacional de Lisboa of the translation Compendio Estadistico ou Resumo dos Elementos de Estadistica [A Treatise on Statistics, or A Resume of the Elements of Statistics] [15] of the 1847 Éléments de Statistique: comprenant les principes généraux de cette science, et un aperçu historique de ses progrès by A. Moreau de Jonnès (then in charge of the bureau of Statistiques Générales of France). It is worth noting that   Jonnès clearly expresses his understanding of the scope and limitations of Statistics:
Statistics is the science of natural, social, and political facts expressed in numbers. Its aim is the deep knowledge of society considered in its nature, elements, economy, situation, and changes. It uses the language of numbers, which is as essential to it as figures are to geometry and symbols are to algebra. It constantly proceeds through numbers, which gives it the precision and certainty characteristic of the exact sciences. Works that bear its name, without having its object or language, do not belong to it, as they fall outside the conditions of its existence. Thus, statistics without numbers, or whose numbers do not express natural, social, and political facts, do not deserve the name they bear. The same can be said of moral and intellectual statistics, because it is a ridiculous claim to want to subject the soul and passions to calculation, and to compute, as if they were defined and comparable units, the movements of the spirit and the phenomena of human intelligence.
For detailed information on Jonnès and his influence, see Paixão Santos (2017) [16].
The politician António José de Ávila (1807-1881) — the presumed inspiration of Eça de Queiroz’s (1925) posthumous satyrical novel O Conde de Abranhos(4)(4)(4)Ávila prohibited a series of liberal conferences (Conferências do Casino), and the outstanding novelist Eça de Queiroz used him as model for the ridiculous politician Conde de Abranhos in a satyrical novel, published only 25 years after the death of Eça de Queiroz. As challenging of the political establishment as Yes, Minister and Yes, Prime Minister. — also reinforced Alexandre Herculano’s proposals in 1854 in a report he prepared for the Minister of Public Works Fontes Pereira de Melo on the works of the Statistical Congress held in Brussels in 1853. He emphasized the importance of establishing uniform bases for national statistical systems, highlighting their importance for the establishment of international comparisons. Cardoso (2012) [17] claims that Ávila “supports the need for a general population census every 10 years, with the collection of basic identification and social status information (profession, education, social and health conditions, etc.). Among the subjects that should be the object of statistical standardization, he gives special attention to the following: territory and property registry; emigration and migratory movements; agricultural production; industrial statistics (production, employment, wages, capital, etc.); commercial statistics; budgets of the working classes; census of the indigent (poverty and social assistance); levels of education; criminality and repression.” Ávila was again the portuguese official delegate at the Statistical Congress held in Berlin in 1863.
There was also a “folkloric” aspect to the use of Statistics: Sexual taboos were also statistically addressed (Santos Cruz, 1841; Azevedo, 1864); and perhaps in an attempt to finance their expenses, some travelers published accounts of their peripatetic wanderings with some pretension of statistical objectivity, such as Portugal a Voo de Pássara(5)(5)(5)  The Flight of a Female Bird in Portugal, as the novelist Camilo Castelo Branco insultingly translated the title Le Portugal à Vol d’Oiseau by Marie Rattazzi, A. Degorce-Cadot, 1880.. In the posthumous publication Londres Maravilhosa [Wonderful London], in the Notes – Excerpts from the notebooks of M. Teixeira-Gomes(6)(6)(6)  Aside being a diplomat and President of the Portuguese Republic, Manuel Teixeira-Gomes was a notorious writer, author namely of the novel Maria Adelaide, published in 1938, describing the attraction for “nymphets” earlier than the famous Lolita of Vladimir Nabokov. Fialho de Almeida was also an important 19th century portuguese novelist, obviously without any serious background in Statistics., he reveals, with a mixture of surprise and irony, that Fialho de Almeida “is preparing for his excursion to Galiza; he plans to make a complete book with statistics and agronomy about Galiza!”
The teaching of elementary Probability Calculus, included in the program of Infinitesimal Calculus, was introduced in 1865 at the Escola Politécnica de Lisboa [Lisbon Polytechnic School], formerly Colégio dos Nobres [Nobleman College] and later, in 1911, Faculdade de Ciências [Faculty of Sciences]. Statistics gradually detached itself from its subordination to Political Economy, and in the last quarter of the 19th century, the first research works linked to Mathematics appeared, authored by Daniel Augusto da Silva (1866, 1870) and Sidónio Paes (1898).
The third posthumous edition of The Doctrine of Chances by Abraham de Moivre (1756) contains applications of Probability Theory to actuarial calculations regarding the computation of annuities. Daniel Augusto da Silva addressed the calculation of annuities in 1866, thus becoming the patron of Portuguese actuaries. It is an interesting investigation but of little statistical interest, the reference to statistics is limited to the excerpt we reproduce.
  • [Uncaptioned image]
  • [Uncaptioned image]
    The values of the amortization taxes given by Statistics and formula (C) for the two Montepio Mutualist Associations that we analyse fall apart, in opposite directions, from the values they would assume in those associations, in case they would be in normal circunstances.
  • [Uncaptioned image]
Daniel Augusto da Silva (1870) also published a comparative study of population movements in Portugal. Detailed information can be found in Martins (2012) [18], with complements in Martins (2020) [19].
  • The Geographia e Estatística Geral de Portugal e Colónias [Geography and General Statistics of Portugal and the Colonies] by Gerardo A. Pery, from 1875, is an interesting synthesis that, without methodological innovations, plays a pioneering role in the association of Geography with Statistics. Population is also a major theme in this work, as far as statistics is concerned.
  • [Uncaptioned image]
  • [Uncaptioned image]
  • From 1871 to 1899 Manuel Nunes Giraldes was the lecturer in Economics and Statistics at the Faculty of Law in Coimbra; in addition to José Frederico Laranjo’s collaboration, M. Emydgio Garcia gave some lectures on the course. Emydgio Garcia published the corresponding notes in 1894, where Statistics was confined to the title of the pamphlet and a set of general comments, which seems to indicate some backlashes in the teaching of Political Economy at the Faculty of Law.
Chapter IV, entitled General Notion of Statistics, pp. 14-17, is merely a set of debatable aphorisms, for instance “Statistics is generally considered the (concrete and descriptive) science that collects, gathers, and records phenomena or facts of observation, expressing them in numbers and geometric signs.
It is not properly a science; it is a necessary aid, an auxiliary, and moreover, an indispensable factor of science.
It is worth mentioning that “mathematical” Statistics had a slow start, mainly because it depends on the language of Probability. The works of John Graunt (1662) [20] and Halley(1693) [21] on mortality are pointed out as the origin of Demography and, indirectly, Actuarial Science, with the referred works of Daniel Augusto da Silva being late descendants.
Other pioneering works in Mathematical Statistics, notably the posthumous publication where Thomas Bayes (1763) [22] used Probability as the language of Statistics for the first time, the great memoir of Laplace’s (1774) [23] on inference, and Gauss’ (1809) [24] research in which the method of least squares is developed, only came to have an impact on statistical production in Portugal at the end of the 19th century.
  • The first doctoral thesis at the Faculty of Mathematics of the University of Coimbra in the field of mathematical statistics, Introdução à Teoria dos Erros das Observações [Introduction to the Theory of Observational Errors], was defended in 1898 by Sidónio Paes. It proposes a rigorous demonstration of the approximation of sums by Gauss’s law, and of the use of “lesser squares”. It explicitly mentions the work of Bayes, Laplace, Lagrange and Gauss, but without precise referencing.
  • [Uncaptioned image]
In fact, in Sidónio Paes thesis, only the books by Poincaré and Bertrand on Probability Calculus, a work by Chauvenet and another by Estienne are identified in footnotes. Works by Encke, Schiaparelli, Stone, Ellis, Tait, Glaisher, Hagen, de Morgan, Ferrero, Bessel and Crofton are mentioned in the text, but without any precise indication.
At the end of the century, Pedro Afonso André de Figueiredo, honoured with the title of Viscount of Wildik, published Notice Statistique sur le Portugal for the Universal Exhibition of 1900.
Chronological List of XIX-th Century Publications in Statistics(7)(7)(7)  In red are the institutions where the rarest works can be found: FLUL – Faculdade de Letras da Universidade de Lisboa; FDUL – Faculdade de Direito da Universidade de Lisboa; FMUL – Faculdade de Medicina da Universidade de Lisboa; IICT – Instituto de Investigação Científica Tropical; IST – Instituto Superior Técnico; ISA – Instituto Superior de Agronomia; FPCEUL – Faculdade de Psicologia e Ciências da Educação da Universidade de Lisboa; FCUL – Faculdade de Ciências da Universidade de Lisboa; ISCSP – Instituto Superior de Ciências Sociais e Políticas; ISEG – Instituto Superior de Economia e Gestão.
Solano Constâncio, F. (1818). Discurso Preliminar [Preliminary Speech], in Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo I, 1–37. (Reproduced in Solano Constâncio, F. (1995). Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, p. 81–99, available at https://www.bportugal.pt/sites/default/ files/ocpep-11.pdf).
Xavier, C. (1820). Considerações sobre a Statistica [Considerations on Statistics], Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo X, p. 134–172. http://books.google.com/books?id=6svQuAvOksoC&printsec=frontcover&dq=editions:04b0YnI0-8bkocSqJH#PRA2-PA134,M1
Balbi, A. (1822). Essai statistique sur le Royaume de Portugal et d’Algarve, comparé aux autres états de l’Europe, et suivi d’un coup’oeil sur l’État actuel des sciences, des lettres et des beaux-arts parmi les portugais des deux hémishéres, Reych et Gravier, libraires, Paris.
https://books.google.fr/books?id=_Jo2AAAAMAAJ&printsec=frontcover#v=onepage&q&f=false (There is a modern edition, Imprensa Nacional – Casa da Moeda, Lisboa, 2004.)
Balbi, A. (1822a). Variétés politiques et statistiques de la monarchie portugaise, Rey et Gravier, Imp. de Casson, Paris. (reedição: Nabu Press, 2010)
(https://books.google.pt/books/about/Vari%C3%A9t%C3%A9s_politico_statistiques_sur_la.html?id =KmrIByr7oocC&redir_esc=y)
Solano Constâncio, F. (1821). An inquiry concerning population, etc…, ou investigação acerca da população e da faculdade de multiplicação da espécie humana; obra destinada a refutar a doutrina do ensaio de M. Malthus sobre esse assunto por W. Godwin, Londres, 1820 [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr.  Malthus’s essay on this subject by W. Godwin, London, 1820], Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo XII, 63–104. (Reproduced in Solano Constâncio, F. (1995). Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, pp. 171–192, available at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf.)
Solano Constâncio, F. (1822). Variétés politico-statistiques sur la monarchie portugaise, por A. Balbi, Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo XVII, 84–103. (Reproduced in Solano Constâncio, F. (1995). Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, pp. 212–222, available at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf.)
Villela da Silva, L. D. (1828). Observações críticas sobre alguns artigos do Ensaio estatístico do reino de Portugal e Algarves, publicado em Paris por Adriano Balbi [Critical observations on some considerations in Variétés politiques et statistiques de la monarchie portugaise, published in Paris by Adran Balbi], Imprensa Régia, Lisboa. FLUL
Solano Constâncio, F. (1836). Novo Diccionario Critico e Etymologico da Lingua Portugueza [New Critical and Etymological Dictionary of the Portuguese Language], Angelo Francisco Carneiro, Paris. (Available at https://books.google.pt/books?id=HM9TAAAAcAAJ&printsec=frontcover&hl=pt-BR&source=gbs_ge_summary_r&cad=0#v=onepage&q&f=false)
Solano Constâncio, F. (1838). Considerações sobre a estatística [Considerations on Statistics], Armazém de Conhecimentos Úteis, nas Artes e Ofícios, ou Colecção de Tratados, Receitas e Invenções de Utilidade Geral [Repository of Useful Knowledge, in Arts and Crafts, or Collection of Treatises, Recipes, and Inventions of General Utility], J.-P. Aillaud, Paris, 1 (1938), 51–64. (Reprinted in Solano Constâncio, F. (1995). Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, p. 271–277, available at https://www.bportugal.pt/sites/ default/files/ocpep-11.pdf.)
Forjaz de Sampaio, A. P. (1841). Primeiros Elementos da Sciencia da Statistica [First Elements of the Science of Statistics], Imprensa de Trovão & Companhia, Coimbra. New edition (together with Elementos de Economia Política [Elements of Political Economy], Coimbra, 1839), entitled Elementos de Economia Política e Estadística [Elements of Political Economy and Statistics], Coimbra, 1845; new edition in 1852; the 1856 edition was entitled Novos Elementos de Economia Política e Estadística [New Elements of Political Economy and Statistics]; in 1858 a new edition in three volumes has been released, entitled Elementos de Economia Política [Elements of Political Economy]. Further infomation in https://www.arqnet.pt/dicionario/sampaioapf.html. Recent edition: Forjaz de Sampaio, A. (1995), Estudos e Elementos de Economia Política [Studies and Elements of Political Economy], Banco de Portugal, Lisboa with and Introduction by Alcino Pedrosa (p. XIII–XXVIII). In Tomo I, available at https://www.bportugal.pt/sites/default/files/ocpep-10_t1.pdf, reprint of Primeiros Elementos da Ciência da Estatística [First Elements of the Science of Statistics], p. 79–112. In Tomo II, available at https://www.bportugal.pt/sites/default/files/ocpep-10_t2.pdf, reprint of the 1974 edition of Elementos de Economia Política e Estadística [Elements of Political Economy and Statistics], with Primeiras Noções da Teoria da Estadística [First Elements of the Science of Statistics] reprinted on pp. 53–60, with notes 291–332 in pp. 373–420.
Santos Cruz, F.  I. (1841). Da prostituição na cidade de Lisboa [On the protitution at Lisbon], Typ. Lisbonense, Lisboa (reprinted in 1984, D. Quixote, Lisboa).
Jonnès, A. M. (1851). Compendio Estadistico ou Resumo dos Elementos de Estadistica [A Treatise on Statistics, or A Resume of the Elements of Statistics], Imprensa Nacional, Lisboa. (2ªed., de 1856, available at https://books.google.tt/books?id=OOoOAAAAQAAJ&printsec=frontcover#v =onepage&q&f=false). ISA
Marreca, A. O. (1854). Parecer e memória sobre a proposta, que apresentou o Sr. Alexandre Herculano: para que a secção de Ciências Económicas e Administrativas redigisse um Projecto de Estatística [A Review and Memoir on Mr. Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Statistics], Academia Real das Ciências, Lisboa. FLUL
Ávila, A. J. (1854). Relatório sobre os Trabalhos do Congresso de Estatística reunido em Bruxelas em 1853 [Report on the Works of the Congress of Statistics held in Brussels in 1853], Imprensa Nacional, Lisboa.
Ferreira, J. D. (1857). Ensaio sobre os Primeiros Elementos da Theoria da Estadistica [Essay on the First Elements of Statistical Theory], Imprensa da Universidade, Coimbra. FDUL
Ávila, A. J. (1864). Relatório sobre os trabalhos do Congresso Internacional de Estatísticas reunido em Berlin em 1863 [Report on the Works of the Congress of Statistics held in Berlin in 1863], Imprensa Nacional, Lisboa. FDUL
Azevedo, F. P. (1864). História da prostituição e policia sanitária no Porto: seguida de um ensaio estatistico dos dous ultimos annos, tabelas comparativas, etc. [History of prostitution and sanitary police at Porto, followed by a statistical essay on the two elapsed years, comparative tables, etc.], F. Gomes da Fonseca, Porto. FMUL
Silva, D. A. (1866). Amortização annual media das pensões nos principaes montepios de sobrevivencia portuguezes [Average yearly amortizations at the foremost survival Montepio Mutualist Associations in Portugal], Jornal de Sciencias Matematicas, Physicas e Naturaes, Tomo I, 175–187. https://www.marinha.pt/pt/a-marinha/historia/personalidades-ilustres/Documents/Daniel%20Aug usto%20da%20Silva/05A_Amortizacao_Annual_Media_Pensoes.pdf
Silva, D. A. (1870). Contribuições para o estudo comparativo do movimento da população em Portugal [Contributions for the comparative study of the population movements in Portugal], Jornal de Sciencias Mathematicas, Physicas e Naturaes, tomo II, VIII, 255–306. FDUL
Pery, G. A. (1875). Geographia e Estatística Geral de Portugal e Colónias [Geography and General Statistics of Portugal and the Colonies], Imprensa Nacional, Lisboa. https://ia800708.us.archive.org/ 19/items/geographiaeesta02perygoog/geographiaeesta02perygoog.pdf IICT
Laranjo, J. F. (1887). Programma para a cadeira de economia política e estatística da Faculdade de Direito da Universidade de Coimbra [Program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra], Imprensa da Universidade, Coimbra. FLUL
Garcia, M. E (1894). Apontamentos de Algumas Prelecções no Curso de Economia Política e Estatística [Notes of Some Magistral Lectures in the Course of Political Economy and Statistics], Typ. de Luiz Cardoso, Coimbra.
Paes, S. (1898). Introdução à Teoria dos Erros das Observações [Introduction to the Theory of Observational Errors]. Imprensa da Universidade, Coimbra. https://am.uc.pt/item/44025
Wildik, Vicomte de (1900). Notice Statistique sur le Portugal, Aillaud, Paris. FLUL
3 20th century — From Rodolpho Guimarães (1904) to 1974
Karl Pearson, William Sealy Gosset (Student), Ronald Fisher, Jerzy Neyman, and Egon Pearson were major figures in the early developements of Mathematical Statistics, namely in what regards inference about parameters assuming normality, discriminant analysis, and experimental design. These concepts were gradually imported, especially through lecture notes (some published as books) by Varennes e Mendonça at ISA (Instituto Superior de Agronomia), and by Leite Pinto and by Bento Murteira at ISCEF (Instituto Superior de Ciências Económicas e Financeiras). On the other hand, the new methods of Statistics were addressed in a series of publications on “methodological issues” by W. L. Stevens (a disciple of Fisher who worked in the Anthropology Institute of Coimbra between 1942 and 1945), and by his patron E. Tamagnini (1937), and by agronomists influenced by Fisher’s school.
Outside the university context, Guimarães (1931) used statistical tests to investigate public education issues, and Amaro Guerreiro (1947)(8)(8)(8)  An interesting analysis of this book can be found at https://www.deepdyve.com/lp/wiley/manual-de-estat-stica-vwDem3M8IE?utm_source=freeShare&utm_medium=link&utm_campaign=freeShare., a senior civil servant at INE (Instituto Nacional de Estatística) included in his book a description of Pearson distributions (inserts between pp. 246 and 247) and in Chapter 10 (Curve Fitting) in the final section addresses the question of “Fitting to Pearson Type I Curve”; it also contains a chapter (8) on Index Numbers and another (9) on Time Series Analysis.
The teaching of Probability at the Polytechnic School and later at the Faculty of Sciences of Lisbon was crucial to making these developments possible.
Probability
Sidónio Paes’s (1898) thesis, presented at the Faculty of Mathematics of Coimbra University, in a way marked the end of the hegemony of the Faculty of Law at the University of Coimbra regarding the university teaching of Statistics in Portugal.
The research, teaching, and dissemination of Probability Calculus effectively enabled the transition from Descriptive Statistics, sufficient for handling official statistics in the 19th century, to the teaching and research of Inferential Statistics.
Since 1865, a chair of Infinitesimal Calculus including the elementary teaching of Probability Calculus had been taught at the Polytechnic School of Lisbon, and in 1911 a quarterly chair of Probability Calculus was created, which in 1932 became an annual chair, taught by Victor Hugo de Lemos and Pedro Braumann (cf. Fontes de Sousa, 2005 [8]). The lecture notes from Victor Hugo de Lemos’s course was edited in 1945 by J. Cabral Madeira.
  • [Uncaptioned image]
  • As far as publications in the area of Probability are concerned, the first noteworthy event was the publication of a booklet intended to the layman, in 1904, volume 223 of the Biblioteca do Povo e das Escolas [Library for the People and the Schools], Noções sobre Cálculo das Probabilidades, Theoria dos Erros e Méthodo dos Mínimos Quadrados [Notions of Probability Calculus, Theory of Errors and the Least Squares Method], by Rodolpho Guimarães.
As Rui Santos (2008a, 2024a, 2024b) [26, 27, 28] comments, Rodolpho Guimarães (9)(9)(9)  Rodolfo Guimarães (1866-1918) was an officer in the Portuguese army (Colonel) and a professor at the Army School of Lisbon (Military Academy). He also served at the Astronomical Observatory of Tapada in Lisbon. Currently, Guimarães is best known for his research in the History of Mathematics, particularly for publishing a catalog of mathematical works by Portuguese authors during the 19th century and for promoting the work of Pedro Nunes (1502–1578), as well as some contributions in the field of Geometry. For detailed information on Rodolfo Guimarães, see Saraiva, L. (1997) [25]. , in the scant 64 pages of this small volume, manages to present an interesting introduction to the topic.
The second notable publication is by José Freire de Sousa Pinto (1855-1911) — the Noções de Cálculo das Probabilidades para o Estabelecimento das Bases da Estatística [Notions of Probability Calculus for Establishing the Bases of Statistics], analysed in Santos (2008b, 2016) [29, 30]. José de Sousa Pinto, a professor at the University of Coimbra, discusses the use of Probability Calculus in modeling random phenomena. In this work, the author begins by explaining the process of constructing knowledge in any science, concluding that the advancement of knowledge results from the combination of observing the phenomenon and theoretical reasoning. Thus, any theory arises from the observation of a phenomenon, initially being an approximation. Then, the phenomenon is observed once more. This new observation serves to test the initial theory, which is confirmed or refined with new knowledge from the new observation, or even altered if the initial theory is found to be erroneous. It is in this successive chain of dual effort, observation and reasoning, that scientific knowledge advances towards a complete understanding of the phenomenon, where the formalized theory will always occur whenever the phenomenon happens.
It should be noted that the analysis of prerequisites for teaching Statistics was a subject of reflections by Leite Pinto (1945), in Noções de Matemática necessárias ao estudo da Estatística [Notions of Mathematics Necessary for the Study of Statistics], and by Amaro Guerreiro, in Em Torno do Ensino da Estatística [About the Teaching of Statistics], 1952-53.
  • [Uncaptioned image]
  • In 1914, Diogo Pacheco de Amorim (1888-1976) defended his doctoral dissertation at the University of Coimbra on Elementos de Cálculo das Probabilidades [Elements of Probability Calculus], the aim of which was to establish rigorous foundations for the theory of Probability. In fact, this objective failed, and another doctoral thesis on Probability, by Manuel Reis (1929), is very critical of Pacheco de Amorim’s thesis.
  • [Uncaptioned image]
Later, in a unique work by Luís de Albuquerque (1959) on the foundations of the Calculus of Probabilities, those dissertations are not even mentioned. Pacheco de Amorim’s thesis has been exhaustively studied by Rui Santos (2008c) [31].
See, as a simple introduction, Pestana and Velosa (2001) [32] and Pestana, D. (1994) [33]. There is an annotated English translation, Mendonça et al. (2007) [34].
In fact, it was only in 1933 that Kolmogorov (1933) [35] made an axiomatic construction of Probability, inspired by Maurice Fréchet’s ideas on the use of Measure Theory.
  • [Uncaptioned image]
  • The Theory of Measurement was tackled by José Joaquim Dionísio (1957) (who, in what regards Probability, published a brief note on Shannon’s entropy in 1959 in Gazeta de Matemática), and became the topic of choice for Pedro Braumann (1969, 1978).
  • [Uncaptioned image]
Pedro Braumann, who was a fellow at Stanford in 1955-56, was also the author of an excellent exposition on the central theme of classical probability theory, the arithmetic of probability laws, [Introduction to the Study of Limits of Sums of Independent Casual Variables] (1958).
Statistics at the University of Coimbra
The teaching of Statistics descending from the links to Political Economy persisted at the Faculty of Law of the University of Coimbra, giving rise to the publication of various texts (Gouveia et al, 1913; Marnoco e Sousa(10)(10)(10)  The incomplete copy from FDUL did not allow us to access some parts of the book; it should be noted that it has a section on “the sceptics of Statistics”, a subject that is currently widely published. In Chapter III (Statistical Laws), sections 59 and 60 are respectively “Law of large numbers” and “Law of small numbers”. Poisson, wanted his result on the convergence of a succession of binomials to Poisson to be known as the law of small numbers, renamed “law of large numbers” Jacques Bernoulli’s “golden theorem” on the convergence of arithmetic means towards the expectation. Only the expression “law of large numbers” persisted., 1913; Silva and Rocha, 1919; Rocha and Monteiro, 1923; Pereira, 1929).
On the other hand, at the Anthropology Institute of the University of Coimbra, Eugénio Tamagnini had an innovative perspective on the contribution of Statistics in Biometrics research. In Revista da Faculdade de Ciências da Universidade de Coimbra, Tamagnini (1937) published an innovative work in Portugal on analysis of variance. Tamagnini also hired W. L. Stevens, a disciple of Ronald Fisher. Xavier da Cunha (1992) [36] comments that “Tamagnini’s interest in the statistical treatment of data obtained from various populations and in comparing them led him to hire a foreign professor of statistics, Professor W. L. Stevens, who collaborated with him on blood groups and also organised free courses in statistics for students of the exact and natural sciences. Stevens had been a brilliant disciple and collaborator of Dr. Fisher, having stayed in Coimbra ‘making us familiar with statistical methods and their most convenient application’, in Tamagnini’s words. This initiative was undoubtedly very useful.
The entry Mathematical Statistics at the University of São Paulo: Mr. W. L. Stevens [37] informs that “[W. L. Stevens] then joined the staff of the Statistical Department, Rothamsted Experimental Station, for a few months on urgent war work prior to leaving for Portugal to take up a lectureship in statistics at Coimbra at the request of the British Council and the Foreign Office. In 1944 he returned to Britain and took a post as statistician with Imperial Chemical Industries, Ltd., Billingham. In 1947 he joined the staff of the Admiralty Statistical Department under Mr. H. L. Seal”, and it can be conjectured that he had something to do with the British intelligence services during the war.
Fonseca (2011) [38] states that “It is in this context of the necessity to improve the statistical treatment of anthropometric, physiological and demographic data, that one understands the arrival of W. L. Stevens, assigned ‘as a statistician, to guide the application work of modern statistical methods to Biological Sciences and organise initiation courses for professors and students of the same methods’. The ‘Elementary Course of Modern Statistical Methods Applicable to Scientific Investigation’, headed by professor W. L. Stevens, functioned between 1942 and 1944, and is linked to a set of published studies in the series ‘Questões de Método’ [‘Methodogical Issues’]. […] As for the years W. L. Stevens spent in Coimbra, the data which allows us to inquire about his teaching activity is frankly scarce. We know that the course he led functioned between 1942 and 1944, but the apparent absence of Stevens’ curriculum vitae in the archives of the University of Coimbra, and, moreover, the inexistence of annuals of the same institution for the period between 1943 and 1947 (encompassing, thus, the years during which the course led by W. L. Stevens was lectured), conditioned our investigation. What was the programme of the course? Who attended it? What bibliography was recommended? Why was it extinct after only two years? These are just some of the questions that (at least) for now, will remain without an answer.”.
Between 1942 and 1945 Stevens published a remarkable set of works which he categorised as Questions of Method — namely on the $\chi^{2}$ significance test, discriminant analysis, statistical distributions, parameters estimation —, disseminating innovations in mathematical statistics by Karl Pearson, Ronald Fisher, Egon Pearson, Jerszy Neyman and their schools.
Stevens also published, with Aureliano Quintanilha and Hugo Ribeiro (1942), an interesting article on the application of Probability Calculus in the investigation of non sexual reproduction by plasmogamy followed by karyogamy in the higher fungi basidiomycetes, in the Gazeta de Matemática (both in the index and in the article the initials W and L are interchanged).
Statistics at the Technical University of Lisbon
Several graduates in Mathematical Sciences, Mathematics and Geographical Engineering from FCUL, where they had studied Probability, Errors and Statistics, became professors at institutions that developed the teaching of Mathematical Statistics, namely at Institutes of the Technical University of Lisbon. Special mention should be made of
  • Manuel Zaluar Nunes (1907-1967), who started as an assistant at ISCEF, before going to study in Paris as a fellow of the Instituto de Alta Cultura, published in 1933 with Mário Santos a concise exposition of Elementos de Cálculo das Probabilidades e de Estatística Matemática [Elements of Probability Calculus and Mathematical Statistics]: written in accordance with the program of the 2nd chair of the Instituto Superior de Ciências Económicas e Financeiras. Founder of the Gazeta de Matemática, Portugaliae Mathematica, and the Sociedade Portuguesa de Matemática [Portuguese Mathematical Society], he was an Assistant Professor at FCUL and later a Full Professor at ISA, a position from which he was dismissed in 1947 for political reasons, having to migrate first to France, where he was a research assistant at the Centre National de la Recherche Scientifique, and later to Brazil, where he was hired as a professor at the University of Recife.
  • Francisco Paulo Pinto Leite (1902-2000), who was a lecturer at the Sorbonne from 1931 to 1933. One of the founders of the Portuguese Mathematical Society, he was appointed Full Professor at ISCEF in 1940, retiring in 1973. There are various versions of the course notes for the Statistics course he taught at ISCEF, and he was also the first lecturer, in 1952-53, of the chair of Econometrics, with Henri Guitton (University of Dijon) and José de Castañeda (University of Madrid), having as assistant Manuel Jacinto Nunes (further information in Machado and Santos Silva, 2002 [39]).
    It should be noted that 1953 saw the publication of Trabalhos do Seminário de Econometria Dirigido pelo Prof. H. O. Wold and the translation of Wold’s book on dynamic systems by Jacinto Nunes, and that Mira Fernandes (who prefaced the book by Mário Santos and Zaluar Nunes) consistently supported research in Statistics and Econometrics, see the series of eight volumes of Estudos de Matemática, Estatística e Econometria [Studies in Mathematics, Statistics and Econometrics] that he edited between 1956 and 1964.
  • José Sebastião e Silva (1914-1972), a fellow of the Instituto de Alta Cultura in Rome from 1942 to 1946, in 1951 won the public competition for the position of Full Professor at the Instituto Superior de Agronomia, taking charge of the chairs of General Mathematics and Calculus of Probabilities. For the latter chair, in 1954-55, he prepared didactic texts, collected in the course notes Cálculo das Probabilidades [Calculus of Probabilities] published at ISA, with an appendix from 1957-58 on correlation and regression, adjustment by the method of least squares, and Student’s $t$ and Fisher-Snedecor’s $F$ statistics, cf. Neves (2005) [40], pp.222–223. In 1991 Calouste Gulbenkian Foundation republished this work, now unfortunately out of print.
Perhaps as a result of the influence of the work developed by Sir Ronald Fisher and his collaborators at Rothamsted, new areas of Statistics were developed at ISA, particularly in the area of Experimental Design (Costa, 1933; Carvalho, 1945, 1946) and Linear Models (Pestana, 1957).
Professor Varennes e Mendonça taught Probabilities at ISA before the appointment of Professor Sebastião e Silva, publishing a brief course on Noções de Cálculo das Probabilidades [Notions of Probability Calculus] (1950), for details cf. Neves (2005) [40], pp. 224–225. Varennes e Mendonça early on realized the difficulties of translating statistical terms, publishing an interesting article on terminology in 1942.
  • [Uncaptioned image]
  • Murteira and Madureira (1958) supplied interesting translations, but it wasn’t until much later that the Specialised Committee on Statistical Nomenclature of the Portuguese Statistical Society and the Brazilian Statistical Association, whose members were Carlos Daniel Paulino (President); Dinis Pestana; João Branco; Lucília Carvalho; Alfredo Egídio dos Reis; Júlio Singer (Chairman); Lúcia Barroso; Wilton Bussab, prepared the Statistics Glossary English-Portuguese, Portuguese Statistical Society / Brazilian Statistical Association https://spestatistica.pt/glossario, which continues to evolve; see also Ventura, C.; Gomes, M.I.; Pestana, D.; Pestana, P. (2013). Glossary of Statistical Terms: German, French, English, Portuguese, INE, Lisbon.
Publishers invest in Statistics — translations and dissemination
The translation, in 1945, of Snedecor’s Statistical Methods Applied to Experiments in Agriculture and Biology certainly had a major impact on the development of Statistics in Portugal. The popularisation of interest in Probability and Statistics, with an increase in the number of students taking courses in these areas, finally aroused the interest of publishers.
  • [Uncaptioned image]
  • Publicações Europa-América pioneered this move to publish statistical books, with the publication of Gustavo de Castro’s books in the collection it Saber [Knowledge] (1959, 1965), an author chosen probably because he had previously published a book on Inferência Estatística [Statistical Inference] at LNEC [National Laboratory of Civil Engineering] in 1953.
  • [Uncaptioned image]
This was followed by the publication of the translation of Tippett’s Statistics (1955), and in 1973-75 Europa-América published the two volumes of P. Dagnelie’s Théorie et Méthodes Statistiques: Applications Agronomiques, translated by A. St. Aubyn, and in 1982 the translation of M. L. Lévy’s Comprendre les Statistiques.
  • [Uncaptioned image]
  • [Uncaptioned image]
  • Editorial Estúdios COR also ventured out with the publication in 1968 of a translation of Gnedenko and Khinchine’s book An Elementary Introduction to the Theory of Probability, with a preface and translation by António Simões Neto; and Despertar published a translation of one of the highest quality and most successful introductory books, Moroney’s Facts from Figures, in 1969.
  • [Uncaptioned image]
Some highlights
The books by Guimarães (1904) and Sousa Pinto (1913) were analysed by Rui Santos, and those by Varennes e Mendonça (1950), Sebastião e Silva (1954) and Pestana (1957) by Manuela Neves [40], as already mentioned.
Of the others, we think the publications by Santos and Zaluar Nunes (1933), Braumann (1958), Reis and Sarmento (1960) and Dias Agudo (1961) deserve special mention.
The publications by E. Tamagnini (1937) and W. L. Stevens (1942-45) in the Revista da Faculdade de Ciências da Universidade de Coimbra [Review of the Faculty of Sciences of Coimbra University] also deserve special mention.
  • The book by Santos and Zaluar Nunes (1933) has a modernity that contrasts sharply with the 1932 book by Bueno y Martins, which has a perspective closer to official statistics.
    • [Uncaptioned image]
    • However, when it comes to descriptive statistics, Bueno y Martins’ book is more detailed, for example, in addition to the arithmetic mean, he considers geometric, harmonic and weighted averages; and in what regards graphical representations, Bueno y Martins discusses cartograms, stereograms, presenting various interesting pictograms, and discusses as well as association and contingency, and correlation.
    • [Uncaptioned image]
    It should be noted, however, that some of the more elaborate graphics in Bueno y Martins (1932) come from other sources. For example, the extra text between pages 138 and 139 is a reproduction of the cover of the August 1930 issue of the Boletim da Direção Geral de Estatística [Bulletin of the General Directorate of Statistics], see the Catálogo Bibliográfico [Bibliographic Catalogue], p.19, published by INE in 2013. Bueno y Martins’ main statistical reference is the monograph by Yule (the bibliographical sources used are indicated at the end of each section, which is very useful), and it is worth noting that Bueno y Martins indicates several books on Probability in the final bibliography, including A Treatise on Probability by Maynard Keynes.
    • [Uncaptioned image]
    • Table of Contents:
      FOREWORD.
      Part I — Theory of Probability.
      Chapter I — General Principles.
      Chapter II — Random variables.
      Chapter III — Fundamental theorems — Probability laws.
      Chapter IV — Probability of causes.
      Part II — Theory of Errors.
      Chapter I — Fundamental Principles.
      Chapter II — Direct observations.
      Chapter III — Indirect observations.
      Chapter IV — Interpolation.
      Part III — Mathematical Statistics.
      Chapter I — Fundamental Problems.
      Chapter II — Schemes.
      Chapter III — Statistical classification.
      Chapter IV — Mean values.
      Chapter V — Correlation.
      Chapter VI — Covariation.
      NOTES.
Appendix 1
Short Courses, Congresses of the Portuguese Statistical Society, downloadable at https:// www.spestatistica.pt/en/publications/category/short-course
Gomes, P. (1998) Tópicos de Sondagens [Survey Topics]
Gomes, M. I.; Figueiredo, F.; Barão, M. I. (1999) 2ªed. 2010, Controlo Estatístico da Qualidade [Statistical Quality Control]
Amaral Turkman, M. A.; Silva, G. (2000) Modelos Lineares Generalizados — da Teoria à Prática [Generalized Linear Models — from Theory to Practice]
Brilhante, M. F.; Pestana, D.; Rocha, J.; Rocha, M. L.; Velosa, S. (2001) 2ª ed. 2011, Inferência Estatística sobre a Localização Usando a Escala [Statistical Inference about Location Using Scale]
Muller, D. (2002) Modelos Heterocedásticos. Aplicações com o software Eviews [Heteroscedastic Models. Applications with Eviews software]
Gonçalves, E.; Mendes Lopes, N. (2003) 2ªed. 2008 Séries Temporais: Modelações Lineares e não Lineares [Time Series: Linear and Non-Linear Modeling]
Branco, J. A. (2004) Uma Introdução à Análise de Clusters [An Introduction to Cluster Analysis]
Braumann, C. (2005) Introdução às Equações Diferenciais Estocásticas e Aplicações [Introduction to Stochastic Differential Equations and Applications]
Rosado, F. (2006) Outliers em Dados Estatísticos [Outliers in Statistical Data]
Pires , A. M.; Branco, J. A. (2007) Introdução aos Métodos Estatísticos Robustos [Introduction to Robust Statistical Methods]
Carvalho, M. L.; Natário, I. C. (2008) Análise de Dados Espaciais [Spatial Data Analysis]
Rocha, C.; Papoila, A. L. (2009) Análise de Sobrevivência [Survival Analysis]
Tenreiro, C. (2010) Uma Introdução à Estimação Não-Paramétrica da Densidade [An Introduction to Non-Parametric Density Estimation]
Cabral, M. S.; Gonçalves, M. H. (2011) Análise de Dados Longitudinais [Longitudinal Data Analysis]
Salgueiro, M. F. (2012) Modelos com Equações Estruturais [Models with Structural Equations]
M. Ivette Gomes, M. I.; Fraga Alves, M. I., Neves, C. (2013) Análise de Valores Extremos : uma Introdução [Extreme Value Analysis: an Introduction]
Amaral Turkman, M. A.; Paulino, C. D. (2015) Estatística Bayesiana Computacional — uma Introdução [Computational Bayesian Statistics — an Introduction]
Brilhante, M. F. (2017) An Introduction to Meta-Analysis [An Introduction to Meta-Analysis]
Amado, C.; Nunes, C.; Sardinha, A. (2019) Análise Estatística de Dados Financeiros [Statistical Analysis of Financial Data]
Afonso, P. M. (2023). Modelação Conjunta de Dados Longitudinais e de Sobrevivência [Joint Modeling of Longitudinal and Survival Data]
Before the SPE Congresses, CEAUL organized conferences, whose proceedings were published:
Actas I Colóquio de Estatística e Investigação Operacional, Lisboa, Centro de Estatística e Aplicações, 1977.
Actas II Colóquio Estatística e Investigação Operacional, Fundão/Covilhã, CEAUL/SPEIO, 1981.
Actas III Colóquio Estatística e Investigação Operacional, CEAUL/SPEIO, Tróia, 1985.
Proceedings of the SPE Congresses:
A Estatística e o Futuro e o Futuro da Estatística [Statistics and the Future and the Future of Statistics], Actas do I Congresso da SPE, 1993, Vimeiro; ed.  Dinis Pestana, Antónia Turkman, João Branco, Luísa Canto e Castro e Ana Pires.
Actas do II Congresso da SPE, 1994, Luso; ed.  M. Nazaré Mendes Lopes, M. Esmeralda Gonçalves, M. Emília Nogueira, Ana Cristina Rosa e Helena Ferreira.
Bom Senso e Sensibilidade: Traves Mestras da Estatística [Common Sense and Sensitivity: Main Supports of Statistics], Actas do III Congresso da SPE, 1995, Guimarães; ed.  João Branco, Paulo Gomes e Jorge Prata.
A Estatística a Decifrar o Mundo [Statistics Deciphering the World], Actas do IV Congresso da SPE, 1996, Funchal; ed.  Rita Vasconcelos, Isabel Fraga Alves, Luisa Canto e Castro e Dinis Pestana.
Estatística – a Diversidade na Unidade [Statistics – Diversity in Unity], Actas do V Congresso Anual da Sociedade Portuguesa de Estatística, 1997, Curia; ed.  Manuela Souto de Miranda e Isabel Pereira.
Afirmar a Estatística: um Desafio para o Século XXI [Affirming Statistics: a Challenge for the 21st Century], Atas do VI Congresso da SPE, 1999, Tomar; ed.  Carlos Daniel Paulino, António Pacheco, Ana Pires e Ferreira da Cunha.
Um Olhar sobre a Estatística [A Look at Statistics], Actas do VII Congresso da SPE, 2001, Ofir; ed.  Pedro Oliveira e Emília Athayde.
A Estatística em Movimento [Statistics in Motion], Actas do VIII Congresso da SPE, 2001, Peniche; ed.  Maria Manuela Neves, Jorge Cadima, Maria João Martins e Fernando Rosado.
Novos Rumos em Estatística [New Directions in Statistics], Actas do IX Congresso da SPE, 2002, Ponta Delgada; ed.  Lucília Carvalho, Fátima Brilhante e Fernando Rosado.
Literacia e Estatística [Literacy and Statistics], Actas do X Congresso da SPE, 2003, Porto; ed.  Paula Brito, Adelaide Figueiredo, Fernanda Sousa, Paulo Teles e Fernando Rosado.
Estatística com Acaso e Necessidade [Statistics with Chance and Necessity], Actas do XI Congresso da SPE, 2004, Faro; ed.  Paulo M. M. Rodrigues, Efigénio da Luz Rebelo e Fernando Rosado.
Estatística Jubilar [Jubilee Statistics], Actas do XII Congresso Anual da SPE, 2005, Évora; ed.  Carlos Braumann, Paulo Infante, Manuela M. Oliveira, Russell Alpízar-Jara e Fernando Rosado.
A Ciência Estatística [Statistical Science], Actas do XIII Congresso da SPE, 2006, Ericeira; ed.  Luísa Canto e Castro, Eugénia Graça Martins, Cristina Rocha, M. Fernanda Oliveira, Margarida Mendes Leal e Fernando Rosado.
Estatística, Ciência Interdisciplinar [Statistics, Interdisciplinary Science], Actas do XIV Congresso da SPE, 2007, Covilhã; ed.  Maria Eugénia Ferrão, Célia Nunes e Carlos Braumann.
Estatística: da Teoria à Prática [Statistics: from Theory to Practice], Actas do XV Congresso da SPE, 2008, Lisboa; ed. Manuela Magalhães Hill, Manuel Alberto Ferreira, José G. Dias, Maria de Fátima Salgueiro, Helena Carvalho, Paula Vicente e Carlos Braumann.
Estatística: Arte de Explicar o Acaso [Statistics: Art of Explaining Chance], Actas do XVI Congresso Anual da SPE, 2009, Vila Real; ed. Irene Oliveira, Elisabete Correia, Fátima Ferreira, Sandra Dias e Carlos Braumann.
Advances in Regression, Survival Analysis, Extreme Values, Markov Processes and Other Statistical Applications, Selected Papers of the Statistical Societies, Springer, XVII Congresso Anual da SPE, 2013, Sesimbra; Ed. João Lita da Silva, Frederico Caeiro, Isabel Natário, Carlos A. Braumann, Manuel L. Esquível and João Tiago Mexia.
Recent Developments in Modeling and Applications in Statistics, Selected Papers of the Statistical Societies, Springer, XVIII Congresso Anual da SPE, 2013, São Pedro do Sul; ed, Paulo Eduardo Oliveira, Maria da Graça Temido, Carla Henriques e Maurizio Vichi.
New Advances in Statistical Modeling and Applications, Selected Papers of the Statistical Societies, Springer, XIX Congresso Anual da SPE, 2014, Nazaré; ed.  António Pacheco, Rui Santos, Maria do Rosário Oliveira e Carlos Daniel Paulino.
Estatística: Novos Desenvolvimentos e Inspirações [Statistics: New Developments and Inspirations], Actas do XX Congresso Anual da SPE, 2013, Porto; ed. Manuela Maia, Pedro Campos e Pedro Duarte Silva.
Estatística: a ciência da incerteza [Statistics: the science of uncertainty], Atas do XXI Congresso da SPE, 2014, Aveiro; ed. Isabel Pereira, Adelaide Freitas, Manuel Scotto, Maria Eduarda Silva e Carlos Daniel Paulino. https://www.spestatistica.pt/publicacoes/publicacao/estatIstica-ciencia-incerteza
Estatística: progressos e aplicações [Statistics: Progresses and Applications], Atas do XXII Congresso da SPE, 2015, Olhão; ed. Clara Cordeiro, Conceição Ribeiro, Carlos Sousa, Maria Helena Gonçalves, Nelson Antunes e Maria Eduarda Silva. https://www.spestatistica.pt/publicacoes/ publicacao/estatIstica-progressos-aplicacoes
Atas do XXIII Congresso da SPE [Proceedings of the XXIII SPE Congress], 2017, Lisboa; ed. Maria de Fátima Salgueiro, Paula Vicente, Teresa Calapez, Catarina Marques e Maria Eduarda Silva. https://www.spestatistica.pt/publicacoes/publicacao/atas-do-xxiii-congresso-da-spe
Estatística: Desafios Transversais às Ciências com Dados [Transversal challenges to Sciences with Data], Atas do XXIV Congresso da Sociedade Portuguesa de Estatística, 2021, Amarante ed. Paula Milheiro, António Pacheco, Bruno de Sousa, Isabel Fraga Alves, Isabel Pereira, Maria João Polidoro e Sandra Ramos. https://www.spestatistica.pt/publicacoes/publicacao/estatistica-desafios-transversais-ciencias-com-dados
Recent Developments in Statistics and Data Science, Atas do XXV Congresso da Sociedade Portuguesa de Estatística, 2021, Évora (ed. Regina Bispo, Lígia Henriques-Rodrigues, Russell Alpizar-Jara e Miguel de Carvalho). https://www.spestatistica.pt/publicacoes/publicacao/recent-developments-statistics-and-data-science
New Frontiers in Statistics and Data Science, Springer Proceedings in Mathematics & Statistics, Springer, XXVI Congresso da SPE, 2024, Guimarães (ed. Lígia Henriques-Rodrigues, Raquel Menezes, Luís Meira Machado, Miguel de Carvalho e Susana Faria), in print.
Appendix 2
Index of the Manual de Estatística Médica (Manual of Medical Statistics) by Carlos Santos Reis and Alexandre Sarmento.
Introduction 3
1 – Fundamental Concepts 6
2 – Statistical Methods 14
   1 – Definition 14
   2 – Objectives 14
   3 – Purpose 14
   4 – Feature 14
   5 – Observation methods 14
   6 – Applications 14
   7 – Limitations 14
First part
Application of the statistical method 25
3 – Planning 26
4 – Organization 28
5 – Preparation 31
   1 – The collection or collection of data 32
      1.1 – Harvest method 33
      1.2 – Observation methods 34
      1.3 – Scope of investigations 36
      1.4 – Investigation Modalities 37
      1.5 – Source of data 42
      1.6 – Preparation of questionnaires 43
      1.7 – Pilot survey 46
      1.8 – Sources of errors in surveys 48
      1.9 – Observation record 52
6 – Data verification 56
   1 – Critique of results 56
   2 – Data ordering 58
   3 – Data classification 61
   4 – Organization of a frequency distribution 64
   5 – Enumeration of individuals 68
   6 – Coding of results 75
7 – Presentation of data 81
   1 – Text presentation 81
   2 – Tabular presentation 82
8 – Graphical presentation 92
   1 – General considerations 92
   2 – Main fundamentals and criticisms of graphic presentation 93
   3 – Rules for constructing graphs 93
   4 – Definition and major divisions in graphic presentation 95
9 – Use 113
   1 – Absolute values 114
   2 – Derived values 114
10 – Description of statistical series 124
   1 – Qualitative characters 125
   2 – Quantitative characters 126
11 – Position measurements 129
   1 – Median 130
   2 – Arithmetic mean 133
   3 – Fashion 137
   4 – Geometric mean 138
   5 – Comparison of main position measurements 142
12 – Dispersion and shape measurements 145
   1 – Dispersion measurements 145
   2 – Shape measurements 152
   3 – Presentation of results 153
13 Simplified calculation of parameters 154
   1 – Fashion 155
   2 – Median 156
   3 – Arithmetic mean 157
   4 – Variance 160
   5 – Joint calculation of mean and variance 164
14 – Interpretation of data. Statistical schemes 167
15 – Some theoretical distributions 174
   1 – Notion of probability and law of large numbers 175
   2 – Binomial distribution 178
   3 – Normal distribution 182
   4 – Poisson Distribution 189
   5 – Variable transformation 191
16 – Sampling 193
   1 – Sampling Deformation 194
   2 – Confidence Limits 196
   3 – Sample representativeness 196
   4 – Sampling Types 197
   5 – Number of observations 202
   6 – Sampling Problems 204
17 – Estimation and confidence limits 205
   1 – Estimation and confidence limits of the mean 207
   2 – Estimation and confidence limits of variance 210
   3 – Estimation and confidence limits of a proportion 210
18 – Compliance Tests 215
   1 – Conformity tests for parameters 219
   2 – Distribution Conformance Tests 220
19 – Homogeneity problem 232
   1 – Comparison of two means 234
   2 – Comparison of two variances or two standard deviations 245
   3 – Comparison of two rates or proportions 248
Appendix
Calculation of factorials 257
20 – Comparison of multiple samples 263
   1 – Simultaneous comparison of multiple means 265
   2 – Simultaneous comparison of several variances 304
   3 – Simultaneous comparison of several proportions 306
21 – Statistical dependence 309
  I – Relationships between two qualitative characters. Notions of independence and association 309
   1 – Presentation of results 310
[…]
   3 – Association Issues 312
   4 – Existence of the association 312
   5 – Independence Tests 314
   6 – Different types of problems 321
   7 – Association Measures 329
22 – Statistical dependence 332
  II – Relationships between two quantitative characters. Understanding correlation and regression 332
   1 – Understanding correlation and regression 333
   2 – Correlation Types 335
   3 – Presentation of results 336
   4 – Correlation form 339
   5 – Regression Line 342
   6 – Regression standard error 350
   7 – Correlation coefficient 351
   8 – Numerical calculation of correlation parameters 353
   9 – Interpretation of correlation coefficients 364
   10 – Significance of a correlation coefficient 366
   11 – Correlation applicability conditions 368
   12 – Advantages of correlation 368
   13 – Comparison of two correlation coefficients 372
   14 – Correlation ratio 375
   15 – Determination and alienation coefficients 376
   16 – Correlation of the number of cases 376
   17 – Position correlation 377
   18 – Conclusions 378
23 – Statistical adjustment 379
   1 – Adjustment concept 380
   2 – Adjustment Methods 380
   3 – Linear adjustment 388
   4 – Curvy Fit 393
   5 – Adjustment to theoretical distributions 400
24 – Checking the normality of a series 406
   1 – Graphic appreciation 407
   2 – Cumulative frequencies 408
   3 – Straight from Henri 409
   4 – Significance of kurtosis and asymmetry measurements 411
   5 – Fitting to a normal distribution 415
[…]
Method    1 – Principles of the 420 method
   2 – Checking the normality of a distribution 424
   3 – Calculation of lethal dose 425
26 – Chronological series 433
   1 – Definition 433
   2 – Movements of a chronological series 434
   3 – Guidance to follow when studying these series 434
   4 – Data preparation 435
   5 – Graphical representation 436
   6 – Analysis of time series 440
   7 – Study of the general trend 441
   8 – Seasonal Trend 444
   9 – Cyclic movements 453
   10 – Comparison of time series 453
27 – Index numbers 456
   1 – Concept 457
   2 – Types of indexes 458
   3 – Conditions for displaying an index number 462
   4 – Problem of creating an index 462
   5 – Random errors in index calculations 464
Second part
Basic notions of biostatistics 465
   0 – Settings 465
   1 – Population 466
      1.1 – Population status 466
      1.2 – Population movement 478
      1.2.1 – Birth 478
      1.2.2 – Mortality 486
      1.2.3 – Demographic balance 497
Third part
Annex I – Statistical terminology 500
Annex II – Statistical symbols 501
Bibliography 503
Appendix 3
Index of Introdução à Estatística (Introduction to Statistics), notes for the course on General Mathematics for the degrees in Biological Sciences, Geological Sciences, and Adjunct Professors, based on the lectures of Professor Dias Agudo, compiled by Moreira Campos / Waldemar Nunes, Lisbon 1961/62
Chapter I – Fundamental Concepts of Probability Theory
      1.1 – Introduction
      1.2 – A fundamental principle
      1.3 – “Arrangements and Combinations”
      1.4 – Event space and probability space
      1.5 – Probabilities of the sum and product of events.
             Conditional probabilities.
             Incompatible and independent events
      1.6 – Problem with repeated tests.
             Binomial formula
Chapter II – Random Variables and Distribution Functions
      2.1 – Casual variables
      2.2 – Distribution function
      2.3 – Classification of distributions
      2.4 – Functions of casual variables
      2.5 – Percentiles of a distribution
Chapter III – Discrete Distributions
      3.1 – Probability function; elementary probabilities
      3.2 – Mean value
      3.3 – Moments
      3.4 – Important examples of discrete distributions
             Poisson distribution
             Binomial Distribution
Chapter IV – Continuous Distributions*
      4.1 – Probability density
      4.2 – Mean value. Moments. Median
      4.3 – Important example of continuous distribution: Normal distribution
             4.3.1 – Reduced normal distribution
             4.3.2 – General normal distribution
Chapter V – Sums of Random Variables
      5.1 – Mean value and variance of the sum
      5.2 – Distribution of the sum of casual variables
      5.3 – Normal distribution as an approximation of binomials
Chapter VI - Sampling
             Data Presentation and Description
      6.1 – Introduction
      6.2 – Frequency distribution; histograms; data classification
      6.3 – Location and dispersion measurements
Chapter VII – Parameter Estimation
      7.1 – Introduction
      7.2 – Mean value estimation
      7.3 – Variance estimation
Chapter VIII – Hypothesis Tests
      8.1 – Introduction
      8.2 – Hypothesis tests on the mean value
      8.3 – Hypothesis tests on the type of distribution
      8.4 – Comparison of two populations

### v2

@@
 How books tell a History of Statistics in Portugal: Works of foreigners, $\boldsymbol{estrangeirados}$, and others
-How Books Tell a History of Statistics in Portugal:
-Works of Foreigners, $\boldsymbol{Estrangeirados}$, and Others
-Dinis Pestana1,3,4, Rui Santos2,3
-1 DEIO, Faculdade de Ciências da Universidade de Lisboa (FCUL)
-2 Escola Superior de Tecnologia e Gestão, Instituto Politécnico de Leiria
-3 Centro de Estatística e Aplicações (CEA/UL)
-4 Instituto de Investigação Científica Bento da Rocha Cabral
+Abstract
+Foreigners and “estrangeirados”, an expression meaning “people going to a foreign country [“estrangeiro”] getting there further education”, had a leading role in the development of Mathematical Statistics in Portugal.
+In what concerns Statistics, “estrangeirados” in the nineteenth century were mainly liberal intellectuals exiled for political reasons.
+From 1930 onwards, the research funding authority sent university professors abroad, and hired foreign researchers to stay in Portuguese institutions, and some of them were instrumental in the importation of new concepts and methods of inferential statistics.
+After 1970, there was a huge program of sending young researchers abroad for doctoral studies.
+At the same time, many new universities and polytechnic institutes have been created in Portugal.
+After that, aside from foreigners who choose to have a research career in those institutions and the “estrangeirados” who had returned and created programs of doctoral studies, others, who hadn’t the opportunity of studying abroad, began to play a decisive role in the development of Statistics in Portugal.
+The publication of handbooks on Probability and Statistics, thesis and core papers in Portuguese scientific journals, and also of works for the layman, reveals how Statistics progressed from descriptive to a mathematical discipline used for inference in all fields of knowledge, from natural sciences to methodology of scientific research.
+keywords:
+History of Statistics, Portugal, 19th Century, 20th Century.
+[…]
+[…]
   • [Uncaptioned image]
@@
 1 Introduction
-Sketching the History of Statistics in Portugal naturally requires discussing the influence of foreigners and “$estrangeirados$”, foreign-influenced individuals(1)(1)(1)  In the 17th and 18th centuries, some Portuguese intellectuals, who were exposed to scientific innovations and the Enlightenment philosophy abroad, had a far-reaching influence on the modernization of culture in Portugal.
+Sketching the History of Statistics in Portugal naturally requires discussing the influence of foreigners and “$estrangeirados$”, foreign-influenced individuals111  In the 17th and 18th centuries, some Portuguese intellectuals, who were exposed to scientific innovations and the Enlightenment philosophy abroad, had a far-reaching influence on the modernization of culture in Portugal.
 Among these “$Estrangeirados$” (as they are referred to in Portuguese history) are Ribeiro Sanches, Luís António Verney, Francisco Xavier de Oliveira (Cavaleiro d’Oliveira), D.
-Luís da Cunha (whose political testament recommended Sebastião José de Carvalho e Melo, later Marquês de Pombal, to King D.
+Luís da Cunha (whose political testament recommended Sebastião José de Carvalho e Melo, later the powerful prime minister Marquês de Pombal, to King D.
 José), and the Marquês de Pombal himself.
@@
 Some of these new estrangeirados, namely Solano Constâncio, Alexandre Herculano, and Oliveira Marreca, were important agents for the dissemination of Statistics in Portugal.
-The politician Passos Manuel, exiled in La Coruña-Plymouth-Belgium-Paris between 1828 and 1832, carried out important reforms during his brief tenure as Minister of the Kingdom in 1836-1837, both in secondary and higher education.
-One of the consequences of his reform was the creation of a course of Economia Política [Political Economy], whose Professor, Adrião Forjaz de Sampaio, included the teaching of Princípios de Statistica [Statistical Primer] in its program from 1841 onwards.
-We use the term $estrangeirado$ in a very general sense: estrangeirados will also be much later Zaluar Nunes (first as a student in Paris, then as a political exile in Brazil, due to the diaspora of intellectuals persecuted by Salazarism), Bento Murteira in London, Gustavo de Castro in Paris, Sebastião e Silva in Italy, Dias Agudo, Tiago de Oliveira, and Pedro Braumann in the USA, and later many scholarship holders who pursued doctorates abroad.
-, without neglecting the role of “others”.
+The politician Passos Manuel, exiled in La Coruña-Plymouth-Belgium-Paris between 1828 and 1832, carried out important education reforms during his brief tenure as Minister of the Kingdom in 1836-1837, both in secondary and higher education.
+One of the consequences of his reform was the creation of a course of Economia Política [Political Economy], whose Professor, Adrião Forjaz de Sampaio, included the teaching of Princípios de Statistica [Statistical Primer] [52] in its program from 1841 onwards.
+We use the term $estrangeirado$ in a very general sense: estrangeirados will also be much later Zaluar Nunes (first as a student in Paris, then as a political exile in Brazil, due to the diaspora of intellectuals persecuted by Salazarism), Bento Murteira in London, Gustavo de Castro in Paris, Sebastião e Silva in Italy, Dias Agudo, Tiago de Oliveira, and Pedro Braumann in the USA, and later many scholarship holders who pursued doctorates abroad., without neglecting the role of “others”.
 Plagiating Mae West, who when asked what kind of men she preferred, candidly replied: “Personally, I like two types of men — domestic and foreign”, in our opinion many “Others”, who hadn’t the opportunity of extended partnerships abroad, had also a relevant role in the development of Mathematical Statistics in Portugal, and their contributions must be recognised and praised as much as those of the estrangeirados.
-Notorious examples are António Simões Neto, a man of culture and one of the early enthusiasts of Bayesian ideas in Portugal, who influenced many students at the Faculty of Sciences of Lisbon (namely Dinis Pestana and Maria Antónia Amaral Turkman), Fernando Rosado (to whom we dedicate this sketch of the history of Statistics in Portugal, thanking his contribution to a “Portugaliae Monumenta Statistica” with his fine edition of the Boletim da Sociedade Portuguesa de Estatística [Bulletin of the Portuguese Statistical Society]), Salomé Cabral or Manuela Neves (recently recipient of the Portuguese Statistical Society Life Achievement Award), whose doctoral dissertations were supervised by Tiago de Oliveira, or Rita Vasconcelos, who contributed so much for the progress of the Probability and Statistics group at the University of Madeira.
+Notorious examples are António Simões Neto, a man of culture and one of the early enthusiasts of Bayesian ideas in Portugal, who influenced many students at the Faculty of Sciences of Lisbon (namely Maria Antónia Amaral Turkman and Dinis Pestana), Fernando Rosado (to whom we dedicate this sketch of the history of Statistics in Portugal, thanking his contribution to a “Portugaliae Monumenta Statistica” with his fine edition of the Boletim da Sociedade Portuguesa de Estatística [148] [Bulletin of the Portuguese Statistical Society]) and of the Memorial da Sociedade Portuguesa de Estatística [147] [Memorial of the Portuguese Statistical Society], Salomé Cabral or Manuela Neves (recently recipient of the Portuguese Statistical Society Life Achievement Award), whose doctoral dissertations were supervised by Tiago de Oliveira, or Rita Vasconcelos, who contributed so much for the progress of the Probability and Statistics group at the University of Madeira.
 Regarding estrangeirados, there is quite a bit of information, but when it comes to the influence of foreigners, the gaps are significant, especially considering the enormous long-distance influence of works already mentioned in the 19th century (Quételet by Forjaz de Sampaio, Bertrand and Poincaré in Sidónio Paes’s thesis) or used as recommended bibliography in courses or research seminars (for example, books by Fisher, Kendall and Stuart, Cramér, Loève, Marek Fisz, Feller, Gnedenko, Cox, Zar, Rohatgi, …).
@@
 During this period, the increasing importance of official statistics and the introduction of Probability teaching at the Polytechnic School contributed to the recognition of Statistics as a mathematical discipline, culminating with Sidónio Paes submission of the first doctoral thesis in the field of Mathematical Statistics in Portugal, at the Faculty of Mathematics of the University of Coimbra.
-Concerning the advancements in Mathematical Statistics in the 20th century up to 1974 (Section 3), we highlight how the teaching of Probability, particularly at the Faculdade de Ciências [Faculty of Sciences], produced excellent professors such as Varennes e Mendonça, Zaluar Nunes, or Sebastião e Silva for Instituto Superior de Agronomia (ISA) [Higher Institute of Agronomy], and Leite Pinto in ISEG.
-These professors, and Bento Murteira, who worked under Maurice Kendall in London, imported the new statistical concepts, particularly those of Pearson (father and son) and Fisher.
+Concerning the advancements in Mathematical Statistics in the 20th century up to 1974 (Section 3), we highlight the consequences of starting the teaching of Probability at the Faculdade de Ciências [Faculty of Sciences].
+Excellent professors developed the teaching of Statistics at the Instituto Superior de Agronomia (ISA) [Higher Institute of Agronomy] — Varennes e Mendonça, Zaluar Nunes, and Sebastião e Silva — and at the Instituto Superior de Economia e Gestão (ISEG) [Higher Institute of Economy and Management], namely Leite Pinto and Bento Murteira, who worked under Maurice Kendall in London.
+These professors imported many new statistical concepts, particularly the innovations created by K.
+Pearson, ‘Student’, Neyman and E.
+Pearson, Fisher, and their co-workers.
 Fisher’s influence was significant at ISA, but we should not overlook the potential role the Institute of Anthropology in Coimbra could have played due to the hiring of Stevens, a disciple of Fisher, who published a series of remarkable “Methodological Issues” in the Revista da Faculdade de Ciências da Universidade de Coimbra [Review of the Faculty of Sciences of the University of Coimbra].
@@
 The Venetian geographer and statistician Adriano Balbi (1782-1848) was in Portugal in 1820, before settling in Paris from 1821 to 1832, where he changed his name to Adrien Balbi.
-It was in Paris, in 1822, that he published the two volumes of the Essai statistique sur le royaume de Portugal et d’Algarve [Statistical Essay on the Kingdom of Portugal and Algarve] and also Variétés politico-statistiques de la monarchie portugaise [Various Political-Statistical Observations on the Portuguese Monarchy].
+It was in Paris, in 1822, that he published the two volumes of the Essai statistique sur le royaume de Portugal et d’Algarve [8] [Statistical Essay on the Kingdom of Portugal and Algarve], and also Variétés politico-statistiques de la monarchie portugaise [9] [Various Political-Statistical Observations on the Portuguese Monarchy].
 Refer to caption Refer to caption Refer to caption
@@
 Figure 1: Covers of Balbi’s books on the Kingdom of Portugal and Villela da Silva’s Critiques
-Some statements in Balbi’s books probably irritated the portuguese establishment, since in 1828 Villela da Silva published negative critical observations on Balbi’s works at the Portuguese Royal Press.
-On the other hand, Balbi’s work was highly praised by Solano Constâncio (1777-1846), who in 1822 published an analysis of Variétés politico-statistiques sur la monarchie portugaise in the Anais das Ciências, Artes, e Letras [Annals of Sciences, Arts, and Letters], Volume XVII, 84–103.(2)(2)(2)  Pages 212–222 of the 1995 reedition of Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], a selection of works by Solano Constâncio [1], published by the Bank of Portugal, edited by José Luís Cardoso, accessible at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf.
+Some statements in Balbi’s books probably irritated the portuguese establishment, since in 1828 Villela da Silva [204] published at the Portuguese Royal Press negative critical observations on Balbi’s works.
+On the other hand, Balbi’s work was highly praised by Solano Constâncio (1777-1846), who in 1822 published an analysis of Variétés politico-statistiques sur la monarchie portugaise [168] in the Anais das Ciências, Artes, e Letras [Annals of Sciences, Arts, and Letters], Volume XVII, 84–103.222  Pages 212–222 of the 1995 reedition of Leituras e Ensaios de Economia Política [171] [Readings and Essays in Political Economy], a selection of works by Solano Constâncio, published by the Bank of Portugal, edited by José Luís Cardoso, accessible at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf.
 The pages cited ahead are from this edition.
-It should be noted, incidentally, that Solano Constâncio, editor of the above mentioned Anais das Ciências, Artes, e Letras, in this critical review refers to the article Considerações sobre a Estatística [Considerations on Statistics] by Cândido Xavier, published in 1820, in Volume X, pp.
+It should be noted, incidentally, that Solano Constâncio, editor of the above mentioned Anais das Ciências, Artes, e Letras, in this critical review refers to the article Considerações sobre a Estatística [Considerations on Statistics] by Cândido Xavier [208], published in 1820, in Volume X, pp.
 134–172.
@@
 From 1822 to 1826 he lived in the United States of America, then returned to Paris, where he died in 1846.
-Machado de Sousa (1979) [2] and Cardoso (1995) [3] are interesting sources of information about the wide range of interests of this remarkable intellectual.
-Regarding Statistics, in the “Preliminary Discourse” of the Anais das Ciências, das Artes e das Letras, Volume I (1818), pp.
+Machado de Sousa [98] and Cardoso [25] are interesting sources of information about the wide range of interests of this remarkable intellectual.
+Regarding Statistics, in the “Preliminary Discourse” [166] of the Anais das Ciências, das Artes e das Letras, Volume I (1818), pp.
 1-37, (pp.
-81-99 of Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Solano Constâncio states that
+81-99 of Leituras e Ensaios de Economia Política [171] [Readings and Essays in Political Economy]), Solano Constâncio states that
 Statistics, a science equally of our times, will also provide material for some articles and for examining the most notable works or news appearing on this subject.” (p. 94)
@@
 Malthus sobre esse assunto por W.
-Godwin, Londres, 1820 [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr.
+Godwin, Londres, 1820 [167] [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr.
 Malthus’s essay on this subject by W.
@@
 The Armazém de Conhecimentos Úteis, nas Artes e Ofícios, ou Colecção de Tratados, Receitas e Invenções de Utilidade Geral [Repository of Useful Knowledge, in Arts and Crafts, or Collection of Treatises, Recipes, and Inventions of General Utility], J.-P.
-Aillaud, Paris, 1838, contains “Considerações sobre a estatística” [“Considerations on Statistics”] (1, 51–64), reproduced on pp.
-271–277 of Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy]; we quote from p.
+Aillaud, Paris, 1838, contains “Considerações sobre a estatística” [170] [“Considerations on Statistics”] (1, 51–64), reproduced on pp.
+271–277 of Leituras e Ensaios de Economia Política[171] [Readings and Essays in Political Economy]; we quote from p.
 276:
 induction from statistical maps will prove the notable utility of this science, which is now indispensable for governing and enlightening the public.
-We find it interesting to note that in the 1821 citation the refutation of Malthus’ exponential population growth, concluding that the human species has very little tendency to increase regularly and permanently foreshadows what Verhulst [4, 5, 6] would later formalize, in 1837, with the logistic model; and that, in the 1838 citation, the notable utility of this science, which is now indispensable for […] enlightening the public has a meaning very close to the famous Thomas Carlyle’s [7] statement in Chap.
+We find it interesting to note that in the 1821 citation [167] the refutation of Malthus’ exponential population growth, concluding that “the human species has very little tendency to increase regularly and permanently” foreshadows what Verhulst [198, 199, 200] would later formalize, in 1837, with the logistic model; and that, in the 1838 citation [170], “the notable utility of this science, which is now indispensable for […] enlightening the public” has a meaning very close to the famous Thomas Carlyle’s [27] statement in Chap.
 II of Chartism: “A judicious man uses statistics, not to get knowledge, but to save himself from having ignorance foisted upon him.”.
@@
 For Solano Constâncio, Statistics has an inferential component that goes beyond its merely descriptive component.
-It is perhaps for this reason that he distinguishes “Estatística”  [Statistics] from “Estadística” [Stadistics], etymogically derived from “Estado” (government), the science of governing or dealing with political affairs, more connected to what is termed official statistics.
-It was Francisco Solano Constâncio who, in 1836, in his Novo Diccionario Critico e Etymologico da Lingua Portugueza [New Critical and Etymological Dictionary of the Portuguese Language], defined, for the first time in Portuguese, Statistics as the science “that deals with the enumeration of everything that forms the strength of a nation”.
+It is perhaps for this reason that he distinguishes “Estatística”  [Statistics] from “Estadística” [Stadistics], etymologically derived from “Estado” (government), the science of governing or dealing with political affairs, more connected to what is termed official statistics.
+It was Francisco Solano Constâncio who, in 1836, in his Novo Diccionario Critico e Etymologico da Lingua Portugueza [169] [New Critical and Etymological Dictionary of the Portuguese Language], defined, for the first time in Portuguese, Statistics as the science “that deals with the enumeration of everything that forms the strength of a nation”.
 He added a list of subjects: population, agriculture, industry, education, public revenues and expenses, military strength, and property distribution.
@@
 Figure 2: Solano Constâncio’ s Novo Diccionario Critico e Etymologico da Lingua Portugueza
-One of the consequences of the 1836 reform by Passos Manuel was the creation of a Chair of Political Economy at the Faculty of Law (resulting from merging the faculties of Laws and of Canons) of the University of Coimbra.
+One of the consequences of the 1836 education reform by Passos Manuel was the creation of a Chair of Political Economy at the Faculty of Law (resulting from merging the faculties of Laws and of Canons) of the University of Coimbra.
 Adrião Forjaz de Sampaio included in the program of the chair an informal presentation of Descriptive Statistics, from the perspective of preparing official statistics, heavily influenced by the journal of works of the Société Française de Statistique Universelle (Fontes de Sousa, 2005 [51]; Ferrão, 2006 [49]).
@@
   • Thus, the pioneer of Statistics teaching at the Faculty of Law of the University of Coimbra was Adrião Forjaz de Sampaio (1810-1874).
-He published Primeiros Elementos da Sciencia da Statistica [First Elements of the Science of Statistics], in which the influence of Quételet is noted.
+He published Primeiros Elementos da Sciencia da Statistica [52] [First Elements of the Science of Statistics], in which the influence of Quételet is noted.
 This manual presents the elements of Descriptive Statistics, giving some prominence to averages and dispersions, with rudimentary mathematical formalism.
 It is interesting to note that in the title, in later editions Forjaz de Sampaio uses “Stadistics” instead of “Statistics”, and that in 1857 José Dias Ferreira (1837-1907), then still a student, published at the University Press an Ensaio sobre os Primeiros Elementos da Theoria da Estadística do Exmo.
-Senhor Adrião Pereira Forjaz de Sampaio [Essay on the First Elements of the Theory of Statistics by the Honorable Adrião Pereira Forjaz de Sampaio].
-In 1887 José Frederico Laranjo published a new program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra, “with the approval of Dr.
+Senhor Adrião Pereira Forjaz de Sampaio [47] [Essay on the First Elements of the Theory of Stadistics by the Honorable Adrião Pereira Forjaz de Sampaio].
+In 1887 José Frederico Laranjo [88] published a new program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra, “with the approval of Dr.
 Manuel Nunes Giraldes”.
-Citing Bastien (1997), [10] “Note that the Chair of Political Economy, after its creation in 1836, and until 1871, was mostly held by Adrião Forjaz de Sampaio and after that date, and until 1899, by Manuel Nunes Giraldes.
+Citing Bastien [15]Note that the Chair of Political Economy, after its creation in 1836, and until 1871, was mostly held by Adrião Forjaz de Sampaio and after that date, and until 1899, by Manuel Nunes Giraldes.
 Despite the strong relationship of Frederico Laranjo with the Chair of Political Economy, particularly in terms of teaching, program development, publication of teaching manuals, and doctoral thesis examinations […] he was never awarded the respective chair.”.
@@
 F.
-Laranjo in Princípios de Economia Política [Principles of Political Economy] (1891), reprinted in [11] do not contain notes on Statistics.
+Laranjo in Princípios de Economia Política (1891) [Principles of Political Economy], reprinted in [89], do not contain notes on Statistics.
 The dichotomy between Estatística/Estadística [Statistics/Stadistics] continued to divide those interested in this new science(s) in the 19th century.
-On March 26, 1852, Alexandre Herculano(3)(3)(3)  A remarkable historian, author of an excellent four volumes very comprehensive História de Portugal [History of Portugal] and of História da Origem e Estabelecimento da Inquisição em Portugal.
+On March 26, 1852, Alexandre Herculano333  A remarkable historian, author of an excellent four volumes very comprehensive História de Portugal [History of Portugal] and of História da Origem e Estabelecimento da Inquisição em Portugal.
 [History of the Origin and Establishment of the Inquisition in Portugal].
 He compiled and edited Portugaliae Monumenta Historica, and was also a highly praised romantic novelist author namely of historical novels.
-(1810-1877) — who was exiled in England and then in France between 1831 and 1832 — proposed to the Royal Academy of Sciences that the section of Economic and Administrative Sciences should “draft the necessary instructions and a series of statistical questions, in harmony with the current state of science, which the Academy should bring to the Government’s attention, seeking to obtain from it an order for civil, ecclesiastical, and military officials, as well as judicial magistrates and elected authorities of any order and denomination, to respond within the scope of their respective actions to the aforementioned questions, with these responses being transmitted to the Academy.”
+(1810-1877) — who was exiled in England and then in France between 1831 and 1832 — proposed to the Lisbon Royal Academy of Sciences that the section of Economic and Administrative Sciences should “draft the necessary instructions and a series of statistical questions, in harmony with the current state of science, which the Academy should bring to the Government’s attention, seeking to obtain from it an order for civil, ecclesiastical, and military officials, as well as judicial magistrates and elected authorities of any order and denomination, to respond within the scope of their respective actions to the aforementioned questions, with these responses being transmitted to the Academy.
 This interest of Herculano in Statistics was not an isolated episode in his vast array of interests and activities.
-Two excerpts from Emigration, in Opúsculos, Volume IV [12] show that Herculano maintained his interest in Statistics, appreciating its contribution to the objective understanding of social phenomena:
+Two excerpts from Emigração [Emigration], in Opúsculos, Volume IV [79] show that Herculano maintained his interest in Statistics, appreciating its contribution to the objective understanding of social phenomena:
 We thus have a set of respectable opinions on the insufficiency of rural wages.
@@
 He was highly influential in the institutionalization of Political Economy in Portugal, for detailed information cf.
-Veríssimo Serrão (1978) [13] and Almodôvar (1995) [14].
+Veríssimo Serrão [201] and Almodôvar [3].
   • [Uncaptioned image]
   • In Parecer e Memoria sobre a proposta, que apresentou o Sr.
-Alexandre Herculano: para que a secção de sciencias económicas e administrativas redigisse um Projecto de Estadística [A Review and Memoir on Mr.
-Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Statistics], Typographia da Academia Real das Sciencias, 1854, Oliveira Marreca returns to the use of the terminology $estadistica$.
+Alexandre Herculano: para que a secção de sciencias económicas e administrativas redigisse um Projecto de Estadística [103] [A Review and Memoir on Mr.
+Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Stadistics], Oliveira Marreca returns to the use of the terminology Estadística.
 Oliveira Marreca was one of the administrators of the Imprensa Nacional de Lisboa [National Press of Lisbon] from 1834 to 1838.
-Arguably Oliveira Marreca influenced the publication at the Imprensa Nacional de Lisboa of the translation Compendio Estadistico ou Resumo dos Elementos de Estadistica [A Treatise on Statistics, or A Resume of the Elements of Statistics] [15] of the 1847 Éléments de Statistique: comprenant les principes généraux de cette science, et un aperçu historique de ses progrès by A.
+Arguably Oliveira Marreca influenced the publication at the Imprensa Nacional de Lisboa of the translation Compendio Estadistico ou Resumo dos Elementos de Estadistica [112] [A Treatise on Statistics, or A Resume of the Elements of Statistics] of the 1847 Éléments de Statistique: comprenant les principes généraux de cette science, et un aperçu historique de ses progrès [111] by A.
 Moreau de Jonnès (then in charge of the bureau of Statistiques Générales of France).
@@
 The same can be said of moral and intellectual statistics, because it is a ridiculous claim to want to subject the soul and passions to calculation, and to compute, as if they were defined and comparable units, the movements of the spirit and the phenomena of human intelligence.
-For detailed information on Jonnès and his influence, see Paixão Santos (2017) [16].
-The politician António José de Ávila (1807-1881) — the presumed inspiration of Eça de Queiroz’s (1925) posthumous satyrical novel O Conde de Abranhos(4)(4)(4)Ávila prohibited a series of liberal conferences (Conferências do Casino), and the outstanding novelist Eça de Queiroz used him as model for the ridiculous politician Conde de Abranhos in a satyrical novel, published only 25 years after the death of Eça de Queiroz.
+For detailed information on Jonnès and his influence, see Paixão Santos [130].
+The politician António José de Ávila (1807-1881) — the presumed inspiration of Eça de Queiroz’s (1925) posthumous satyrical novel O Conde de Abranhos444Ávila prohibited a series of liberal conferences (Conferências do Casino), and the outstanding novelist Eça de Queiroz used him as model for the ridiculous politician Conde de Abranhos in a satyrical novel, published only 25 years after the death of Eça de Queiroz.
 As challenging of the political establishment as Yes, Minister and Yes, Prime Minister.
@@
 He emphasized the importance of establishing uniform bases for national statistical systems, highlighting their importance for the establishment of international comparisons.
-Cardoso (2012) [17] claims that Ávila “supports the need for a general population census every 10 years, with the collection of basic identification and social status information (profession, education, social and health conditions, etc.).
-Among the subjects that should be the object of statistical standardization, he gives special attention to the following: territory and property registry; emigration and migratory movements; agricultural production; industrial statistics (production, employment, wages, capital, etc.); commercial statistics; budgets of the working classes; census of the indigent (poverty and social assistance); levels of education; criminality and repression.” Ávila was again the portuguese official delegate at the Statistical Congress held in Berlin in 1863.
-There was also a “folkloric” aspect to the use of Statistics: Sexual taboos were also statistically addressed (Santos Cruz, 1841; Azevedo, 1864); and perhaps in an attempt to finance their expenses, some travelers published accounts of their peripatetic wanderings with some pretension of statistical objectivity, such as Portugal a Voo de Pássara(5)(5)(5)  The Flight of a Female Bird in Portugal, as the novelist Camilo Castelo Branco insultingly translated the title Le Portugal à Vol d’Oiseau by Marie Rattazzi, A.
-Degorce-Cadot, 1880..
+Cardoso [26] claims that Ávila “supports the need for a general population census every 10 years, with the collection of basic identification and social status information (profession, education, social and health conditions, etc.).
+Among the subjects that should be the object of statistical standardization, he gives special attention to the following: territory and property registry; emigration and migratory movements; agricultural production; industrial statistics (production, employment, wages, capital, etc.); commercial statistics; budgets of the working classes; census of the indigent (poverty and social assistance); levels of education; criminality and repression.” Ávila was again the portuguese official delegate at the Statistical Congress held in Berlin in 1863 [6].
+There was also a “folkloric” aspect to the use of Statistics: Sexual taboos were also statistically addressed (Santos Cruz [156]; Azevedo [7]); and perhaps in an attempt to finance their expenses, some travellers published accounts of their peripatetic wanderings with some pretension of statistical objectivity, such as Portugal a Voo de Pássara [142]555  As the novelist Camilo Castelo Branco insultingly translated the title Le Portugal à Vol d’Oiseau [The Flight of a Female Bird in Portugal,] by Marie Rattazzi, A.
+Degorce-Cadot, 1880.
+“Pássara” means literally “female bird”, but its slang meaning is “cunning”, “smart-ass”, “shrewed-ass”, or even worse..
 In the posthumous publication Londres Maravilhosa [Wonderful London], in the Notes – Excerpts from the notebooks of M.
-Teixeira-Gomes(6)(6)(6)  Aside being a diplomat and President of the Portuguese Republic, Manuel Teixeira-Gomes was a notorious writer, author namely of the novel Maria Adelaide, published in 1938, describing the attraction for “nymphets” earlier than the famous Lolita of Vladimir Nabokov.
-Fialho de Almeida was also an important 19th century portuguese novelist, obviously without any serious background in Statistics., he reveals, with a mixture of surprise and irony, that Fialho de Almeida “is preparing for his excursion to Galiza; he plans to make a complete book with statistics and agronomy about Galiza!”
+Teixeira-Gomes666  Aside being a diplomat and President of the Portuguese Republic, Manuel Teixeira-Gomes was a notorious writer, author namely of the novel Maria Adelaide, published in 1938, describing the attraction for “nymphets” earlier than the famous Lolita of Vladimir Nabokov.
+Fialho de Almeida was also an important 19th century portuguese novelist, obviously without any serious background in Statistics.
+In 1905 he published Cadernos de Viagem.
+Galiza [Travel Notebooks.
+Galiza]., he reveals, with a mixture of surprise and irony, that the novelist and journalist Fialho de Almeida “is preparing for his excursion to Galiza; he plans to make a complete book with statistics and agronomy about Galiza!
 The teaching of elementary Probability Calculus, included in the program of Infinitesimal Calculus, was introduced in 1865 at the Escola Politécnica de Lisboa [Lisbon Polytechnic School], formerly Colégio dos Nobres [Nobleman College] and later, in 1911, Faculdade de Ciências [Faculty of Sciences].
-Statistics gradually detached itself from its subordination to Political Economy, and in the last quarter of the 19th century, the first research works linked to Mathematics appeared, authored by Daniel Augusto da Silva (1866, 1870) and Sidónio Paes (1898).
-The third posthumous edition of The Doctrine of Chances by Abraham de Moivre (1756) contains applications of Probability Theory to actuarial calculations regarding the computation of annuities.
+Statistics gradually detached itself from its subordination to Political Economy, and in the last quarter of the 19th century, the first research works linked to Mathematics were published in 1866 and 1870 (Daniel Augusto da Silva [160, 161]) and later, in 1898, the first doctoral thesis on Mathematical Statistics (Sidónio Paes [129]).
+The third posthumous edition (1756) of The Doctrine of Chances by Abraham de Moivre [40] contains applications of Probability Theory to actuarial calculations regarding the computation of annuities.
 Daniel Augusto da Silva addressed the calculation of annuities in 1866, thus becoming the patron of Portuguese actuaries.
@@
   • [Uncaptioned image]
-    The values of the amortization taxes given by Statistics and formula (C) for the two Montepio Mutualist Associations that we analyse fall apart, in opposite directions, from the values they would assume in those associations, in case they would be in normal circunstances.
+    The values of the amortization taxes given by Statistics and formula (C) for the two Montepio Mutualist Associations that we analyse fall apart, in opposite directions, from the values they would assume in those associations, in case they would be in normal circumstances.
   • [Uncaptioned image]
-Daniel Augusto da Silva (1870) also published a comparative study of population movements in Portugal.
-Detailed information can be found in Martins (2012) [18], with complements in Martins (2020) [19].
-  • The Geographia e Estatística Geral de Portugal e Colónias [Geography and General Statistics of Portugal and the Colonies] by Gerardo A.
+Daniel Augusto da Silva also published a comparative study of population movements in Portugal [161]. Detailed information can be found in Martins [105], with further complements in Martins [106].
+  • The Geographia e Estatística Geral de Portugal e Colónias [136] [Geography and General Statistics of Portugal and the Colonies] by Gerardo A.
 Pery, from 1875, is an interesting synthesis that, without methodological innovations, plays a pioneering role in the association of Geography with Statistics.
@@
 Emydgio Garcia gave some lectures on the course.
-Emydgio Garcia published the corresponding notes in 1894, where Statistics was confined to the title of the pamphlet and a set of general comments, which seems to indicate some backlashes in the teaching of Political Economy at the Faculty of Law.
+Emydgio Garcia published the corresponding notes in 1894 [55], where Statistics was confined to the title of the pamphlet and a set of general comments, which seems to indicate some backlashes in the teaching of Political Economy at the Faculty of Law.
 Chapter IV, entitled General Notion of Statistics, pp.
@@
 The works of John Graunt (1662) [69] and Halley(1693) [78] on mortality are pointed out as the origin of Demography and, indirectly, Actuarial Science, with the referred works of Daniel Augusto da Silva being late descendants.
-Other pioneering works in Mathematical Statistics, notably the posthumous publication where Thomas Bayes (1763) [22] used Probability as the language of Statistics for the first time, the great memoir of Laplace’s (1774) [23] on inference, and Gauss’ (1809) [24] research in which the method of least squares is developed, only came to have an impact on statistical production in Portugal at the end of the 19th century.
-  • The first doctoral thesis at the Faculty of Mathematics of the University of Coimbra in the field of mathematical statistics, Introdução à Teoria dos Erros das Observações [Introduction to the Theory of Observational Errors], was defended in 1898 by Sidónio Paes.
+Other pioneering works in Mathematical Statistics, notably the posthumous publication where Thomas Bayes (1763) [16] used Probability as the language of Statistics for the first time, the great memoir of Laplace’s (1774) [87] on inference, and Gauss’ research (1809) [56] in which the method of least squares is developed, only came to have an impact on statistical production in Portugal at the end of the 19th century.
+  • The first doctoral thesis at the Faculty of Mathematics of the University of Coimbra in the field of Mathematical Statistics, Introdução à Teoria dos Erros das Observações [129] [Introduction to the Theory of Observational Errors], was defended in 1898 by Sidónio Paes.
 It proposes a rigorous demonstration of the approximation of sums by Gauss’s law, and of the use of “lesser squares”.
-It explicitly mentions the work of Bayes, Laplace, Lagrange and Gauss, but without precise referencing.
+It explicitly mentions the works of Bayes, Laplace, Lagrange and Gauss, but without precise referencing.
   • [Uncaptioned image]
@@
 Works by Encke, Schiaparelli, Stone, Ellis, Tait, Glaisher, Hagen, de Morgan, Ferrero, Bessel and Crofton are mentioned in the text, but without any precise indication.
-At the end of the century, Pedro Afonso André de Figueiredo, honoured with the title of Viscount of Wildik, published Notice Statistique sur le Portugal for the Universal Exhibition of 1900.
-Chronological List of XIX-th Century Publications in Statistics(7)(7)(7)  In red are the institutions where the rarest works can be found: FLUL – Faculdade de Letras da Universidade de Lisboa; FDUL – Faculdade de Direito da Universidade de Lisboa; FMUL – Faculdade de Medicina da Universidade de Lisboa; IICT – Instituto de Investigação Científica Tropical; IST – Instituto Superior Técnico; ISA – Instituto Superior de Agronomia; FPCEUL – Faculdade de Psicologia e Ciências da Educação da Universidade de Lisboa; FCUL – Faculdade de Ciências da Universidade de Lisboa; ISCSP – Instituto Superior de Ciências Sociais e Políticas; ISEG – Instituto Superior de Economia e Gestão.
-Solano Constâncio, F.
-(1818).
-Discurso Preliminar [Preliminary Speech], in Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo I, 1–37.
-(Reproduced in Solano Constâncio, F.
-(1995).
-Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, p.
-81–99, available at https://www.bportugal.pt/sites/default/ files/ocpep-11.pdf).
-Xavier, C.
-(1820).
-Considerações sobre a Statistica [Considerations on Statistics], Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo X, p.
-134–172.
-http://books.google.com/books?id=6svQuAvOksoC&printsec=frontcover&dq=editions:04b0YnI0-8bkocSqJH#PRA2-PA134,M1
-Balbi, A.
-(1822).
-Essai statistique sur le Royaume de Portugal et d’Algarve, comparé aux autres états de l’Europe, et suivi d’un coup’oeil sur l’État actuel des sciences, des lettres et des beaux-arts parmi les portugais des deux hémishéres, Reych et Gravier, libraires, Paris.
-https://books.google.fr/books?id=_Jo2AAAAMAAJ&printsec=frontcover#v=onepage&q&f=false (There is a modern edition, Imprensa Nacional – Casa da Moeda, Lisboa, 2004.)
-Balbi, A. (1822a). Variétés politiques et statistiques de la monarchie portugaise, Rey et Gravier, Imp. de Casson, Paris. (reedição: Nabu Press, 2010)
-(https://books.google.pt/books/about/Vari%C3%A9t%C3%A9s_politico_statistiques_sur_la.html?id =KmrIByr7oocC&redir_esc=y)
-Solano Constâncio, F.
-(1821).
-An inquiry concerning population, etc…, ou investigação acerca da população e da faculdade de multiplicação da espécie humana; obra destinada a refutar a doutrina do ensaio de M.
-Malthus sobre esse assunto por W.
-Godwin, Londres, 1820 [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr.
+At the end of the century, Pedro Afonso André de Figueiredo, honoured with the title of Viscount of Wildik, published Notice Statistique sur le Portugal [203] for the Universal Exhibition of 1900.
+Chronology
+(English translation of Portuguese titles inside square brackets.
+Indication of libraries where rare books were found inside angle brackets: $<FLUL>$ – Faculdade de Letras da Universidade de Lisboa; $<FDUL>$ – Faculdade de Direito da Universidade de Lisboa; $<FMUL>$ – Faculdade de Medicina da Universidade de Lisboa; $<IICT>$ – Instituto de Investigação Científica Tropical; $<IST>$ – Instituto Superior Técnico; $<ISA>$ – Instituto Superior de Agronomia; $<FPCEUL>$ – Faculdade de Psicologia e Ciências da Educação da Universidade de Lisboa; $<FCUL>$ – Faculdade de Ciências da Universidade de Lisboa; $<ISCSP>$ – Instituto Superior de Ciências Sociais e Políticas; $<ISEG>$ – Instituto Superior de Economia e Gestão.)
+1818 — F. Solano Constâncio’s [Preliminary Speech] [166].
+1820 — C. Xavier’s [Considerations on Statistics] [207].
+1821 — F.
+Solano Constâncio’s [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr.
 Malthus’s essay on this subject by W.
-Godwin, London, 1820], Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo XII, 63–104.
-(Reproduced in Solano Constâncio, F.
-(1995).
-Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, pp.
-171–192, available at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf.)
-Solano Constâncio, F.
-(1822).
-Variétés politico-statistiques sur la monarchie portugaise, por A.
-Balbi, Anais das Ciências, das Artes e das Letras [Annals of Sciences, Arts, and Letters], Tomo XVII, 84–103.
-(Reproduced in Solano Constâncio, F.
-(1995).
-Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, pp.
-212–222, available at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf.)
-Villela da Silva, L.
-D.
-(1828).
-Observações críticas sobre alguns artigos do Ensaio estatístico do reino de Portugal e Algarves, publicado em Paris por Adriano Balbi [Critical observations on some considerations in Variétés politiques et statistiques de la monarchie portugaise, published in Paris by Adran Balbi], Imprensa Régia, Lisboa.
-FLUL
-Solano Constâncio, F.
-(1836).
-Novo Diccionario Critico e Etymologico da Lingua Portugueza [New Critical and Etymological Dictionary of the Portuguese Language], Angelo Francisco Carneiro, Paris.
-(Available at https://books.google.pt/books?id=HM9TAAAAcAAJ&printsec=frontcover&hl=pt-BR&source=gbs_ge_summary_r&cad=0#v=onepage&q&f=false)
-Solano Constâncio, F.
-(1838).
-Considerações sobre a estatística [Considerations on Statistics], Armazém de Conhecimentos Úteis, nas Artes e Ofícios, ou Colecção de Tratados, Receitas e Invenções de Utilidade Geral [Repository of Useful Knowledge, in Arts and Crafts, or Collection of Treatises, Recipes, and Inventions of General Utility], J.-P.
-Aillaud, Paris, 1 (1938), 51–64.
-(Reprinted in Solano Constâncio, F.
-(1995).
-Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Banco de Portugal, Lisboa, p.
-271–277, available at https://www.bportugal.pt/sites/ default/files/ocpep-11.pdf.)
-Forjaz de Sampaio, A.
-P.
-(1841).
-Primeiros Elementos da Sciencia da Statistica [First Elements of the Science of Statistics], Imprensa de Trovão & Companhia, Coimbra.
-New edition (together with Elementos de Economia Política [Elements of Political Economy], Coimbra, 1839), entitled Elementos de Economia Política e Estadística [Elements of Political Economy and Statistics], Coimbra, 1845; new edition in 1852; the 1856 edition was entitled Novos Elementos de Economia Política e Estadística [New Elements of Political Economy and Statistics]; in 1858 a new edition in three volumes has been released, entitled Elementos de Economia Política [Elements of Political Economy].
-Further infomation in https://www.arqnet.pt/dicionario/sampaioapf.html.
-Recent edition: Forjaz de Sampaio, A.
-(1995), Estudos e Elementos de Economia Política [Studies and Elements of Political Economy], Banco de Portugal, Lisboa with and Introduction by Alcino Pedrosa (p.
-XIII–XXVIII).
-In Tomo I, available at https://www.bportugal.pt/sites/default/files/ocpep-10_t1.pdf, reprint of Primeiros Elementos da Ciência da Estatística [First Elements of the Science of Statistics], p.
-79–112.
-In Tomo II, available at https://www.bportugal.pt/sites/default/files/ocpep-10_t2.pdf, reprint of the 1974 edition of Elementos de Economia Política e Estadística [Elements of Political Economy and Statistics], with Primeiras Noções da Teoria da Estadística [First Elements of the Science of Statistics] reprinted on pp.
-53–60, with notes 291–332 in pp.
-373–420.
-Santos Cruz, F.  I. (1841). Da prostituição na cidade de Lisboa [On the protitution at Lisbon], Typ. Lisbonense, Lisboa (reprinted in 1984, D. Quixote, Lisboa).
-Jonnès, A.
-M.
-(1851).
-Compendio Estadistico ou Resumo dos Elementos de Estadistica [A Treatise on Statistics, or A Resume of the Elements of Statistics], Imprensa Nacional, Lisboa.
-(2ªed., de 1856, available at https://books.google.tt/books?id=OOoOAAAAQAAJ&printsec=frontcover#v =onepage&q&f=false).
-ISA
-Marreca, A.
-O.
-(1854).
-Parecer e memória sobre a proposta, que apresentou o Sr.
-Alexandre Herculano: para que a secção de Ciências Económicas e Administrativas redigisse um Projecto de Estatística [A Review and Memoir on Mr.
-Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Statistics], Academia Real das Ciências, Lisboa.
-FLUL
-Ávila, A.
-J.
-(1854).
-Relatório sobre os Trabalhos do Congresso de Estatística reunido em Bruxelas em 1853 [Report on the Works of the Congress of Statistics held in Brussels in 1853], Imprensa Nacional, Lisboa.
-Ferreira, J. D. (1857). Ensaio sobre os Primeiros Elementos da Theoria da Estadistica [Essay on the First Elements of Statistical Theory], Imprensa da Universidade, Coimbra. FDUL
-Ávila, A.
-J.
-(1864).
-Relatório sobre os trabalhos do Congresso Internacional de Estatísticas reunido em Berlin em 1863 [Report on the Works of the Congress of Statistics held in Berlin in 1863], Imprensa Nacional, Lisboa.
-FDUL
-Azevedo, F.
-P.
-(1864).
-História da prostituição e policia sanitária no Porto: seguida de um ensaio estatistico dos dous ultimos annos, tabelas comparativas, etc.
-[History of prostitution and sanitary police at Porto, followed by a statistical essay on the two elapsed years, comparative tables, etc.], F.
-Gomes da Fonseca, Porto.
-FMUL
-Silva, D.
-A.
-(1866).
-Amortização annual media das pensões nos principaes montepios de sobrevivencia portuguezes [Average yearly amortizations at the foremost survival Montepio Mutualist Associations in Portugal], Jornal de Sciencias Matematicas, Physicas e Naturaes, Tomo I, 175–187.
-https://www.marinha.pt/pt/a-marinha/historia/personalidades-ilustres/Documents/Daniel%20Aug usto%20da%20Silva/05A_Amortizacao_Annual_Media_Pensoes.pdf
-Silva, D.
-A.
-(1870).
-Contribuições para o estudo comparativo do movimento da população em Portugal [Contributions for the comparative study of the population movements in Portugal], Jornal de Sciencias Mathematicas, Physicas e Naturaes, tomo II, VIII, 255–306.
-FDUL
-Pery, G.
-A.
-(1875).
-Geographia e Estatística Geral de Portugal e Colónias [Geography and General Statistics of Portugal and the Colonies], Imprensa Nacional, Lisboa.
-https://ia800708.us.archive.org/ 19/items/geographiaeesta02perygoog/geographiaeesta02perygoog.pdf IICT
-Laranjo, J.
-F.
-(1887).
-Programma para a cadeira de economia política e estatística da Faculdade de Direito da Universidade de Coimbra [Program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra], Imprensa da Universidade, Coimbra.
-FLUL
-Garcia, M.
-E (1894).
-Apontamentos de Algumas Prelecções no Curso de Economia Política e Estatística [Notes of Some Magistral Lectures in the Course of Political Economy and Statistics], Typ.
-de Luiz Cardoso, Coimbra.
-Paes, S. (1898). Introdução à Teoria dos Erros das Observações [Introduction to the Theory of Observational Errors]. Imprensa da Universidade, Coimbra. https://am.uc.pt/item/44025
-Wildik, Vicomte de (1900). Notice Statistique sur le Portugal, Aillaud, Paris. FLUL
+Godwin, London, 1820], [167].
+1822 — A.
+Balbi’s Essai statistique sur le Royaume de Portugal et d’Algarve, comparé aux autres états de l’Europe, et suivi d’un coup’oeil sur l’État actuel des sciences, des lettres et des beaux-arts parmi les portugais des deux hémishéres [8].
+— A. Balbi’s Variétés politiques et statistiques de la monarchie portugaise [9].
+— F. Solano Constâncio’s [Variétés politico-statistiques sur la monarchie portugaise, by A. Balbi] [168].
+1828 — L. D. Villela da Silva’s [Critical observations on some considerations in Variétés politiques et statistiques de la monarchie portugaise, published in Paris by Adran Balbi] [204]. $<FLUL>$
+1836 — F. Solano Constâncio’s [New Critical and Etymological Dictionary of the Portuguese Language] [169].
+— Passos Manuel’s reform of the University, namely creating of a Chair of Political Economy at the Faculty of Law of the University of Coimbra.
+1838 — F. Solano Constâncio’s [Considerations on Statistics] [170].
+1841 — A. P. Forjaz de Sampaio’s [First Elements of the Science of Statistics] [52].
+— F. I. Santos Cruz’s [On the Prostitution at Lisbon] [156].
+1854 — A.
+Oliveira Marreca’s [A Review and Memoir on Mr.
+Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Stadistics] [103].
+$<FLUL>$
+— A. J. d’Ávila’s [Report on the Works of the Congress of Statistics held in Brussels in 1853] [5].
+1856 — A.
+Moreau de Jonnès’ Compendio Estadistico ou Resumo dos Elementos de Estadistica [112], translation of Éléments de Statistique: comprenant les principes généraux de cette science, et un aperçu historique de ses progrès [111].
+$<ISA>$
+1857 — J. Dias Ferreira’s [Essay on the First Elements of Stadistics Theory] [47]. $<FDUL>$
+1864 — creation of the Direção-Geral de Estatística [General Direction of Statistics].
+— A. J. d’Ávila’s [Report on the Works of the Congress of Statistics held in Berlin in 1863], [6]. $<FDUL>$
+— F. P.  Azevedo’s [History of prostitution and sanitary police at Porto, followed by a statistical essay on the two elapsed years, comparative tables, etc.] [7]. $<FMUL>$
+1865 — The teaching of elementary Probability Calculus was included in the program of Infinitesimal Calculus, at the Escola Politécnica de Lisboa [Lisbon Polytechnic School].
+1866 — D. A. da Silva’s [Average yearly amortizations at the foremost survival Montepio Mutualist Associations in Portugal][160].
+1870 — D. A. da Silva’s [Contributions for the comparative study of the population movements in Portugal][161]. $<FDUL>$
+1875 — G. A. Pery’s [Geography and General Statistics of Portugal and the Colonies] [136]. $<IICT>$
+1887 — J. F. Laranjo’s [Program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra] [88]. $<FLUL>$
+1894 — M. E. Garcia’s [Notes of Some Magistral Lectures in the Course of Political Economy and Statistics] [55].
+1898 — Sidónio Paes’ [Introduction to the Theory of Observational Errorss] [129].
+1900 — Vicomte de Wildik’s [Notice Statistique sur le Portugal]. [203] $<FLUL>$
 3 20th Century — From Rodolpho Guimarães (1904) to 1974
-Karl Pearson, William Sealy Gosset (Student), Ronald Fisher, Jerzy Neyman, and Egon Pearson were major figures in the early developements of Mathematical Statistics, namely in what regards inference about parameters assuming normality, discriminant analysis, and experimental design.
+Karl Pearson, William Sealy Gosset (Student), Ronald Fisher, Jerzy Neyman, and Egon Pearson were major figures in the early developments of Mathematical Statistics, namely in what regards inference about parameters assuming normality, discriminant analysis, and experimental design.
 These concepts were gradually imported, especially through lecture notes (some published as books) by Varennes e Mendonça at ISA (Instituto Superior de Agronomia), and by Leite Pinto and by Bento Murteira at ISCEF (Instituto Superior de Ciências Económicas e Financeiras).
@@
 Stevens (a disciple of Fisher who worked in the Anthropology Institute of Coimbra between 1942 and 1945), and by his patron E.
-Tamagnini (1937), and by agronomists influenced by Fisher’s school.
-Outside the university context, Guimarães (1931) used statistical tests to investigate public education issues, and Amaro Guerreiro (1947)(8)(8)(8)  An interesting analysis of this book can be found at https://www.deepdyve.com/lp/wiley/manual-de-estat-stica-vwDem3M8IE?utm_source=freeShare&utm_medium=link&utm_campaign=freeShare., a senior civil servant at INE (Instituto Nacional de Estatística) included in his book a description of Pearson distributions (inserts between pp.
+Tamagnini, and by agronomists influenced by Fisher’s school.
+Outside the university context, Guimarães [74] used statistical tests to investigate public education issues, and Amaro Guerreiro [72]777  An interesting analysis of this book can be found at https://www.deepdyve.com/lp/wiley/manual-de-estat-stica-vwDem3M8IE?utm_source=freeShare&utm_medium=link&utm_campaign=freeShare., a senior civil servant at INE (Instituto Nacional de Estatística) included in his book a description of Pearson distributions (inserts between pp.
 246 and 247) and in Chapter 10 (Curve Fitting) in the final section addresses the question of “Fitting to Pearson Type I Curve”; it also contains a chapter (8) on Index Numbers and another (9) on Time Series Analysis.
@@
 Probability
-Sidónio Paes’s (1898) thesis, presented at the Faculty of Mathematics of Coimbra University, in a way marked the end of the hegemony of the Faculty of Law at the University of Coimbra regarding the university teaching of Statistics in Portugal.
+Sidónio Paes’s 1898 thesis [129], presented at the Faculty of Mathematics of Coimbra University, in a way marked the end of the hegemony of the Faculty of Law at the University of Coimbra in what concerns the university teaching of Statistics in Portugal.
 The research, teaching, and dissemination of Probability Calculus effectively enabled the transition from Descriptive Statistics, sufficient for handling official statistics in the 19th century, to the teaching and research of Inferential Statistics.
-Since 1865, a chair of Infinitesimal Calculus including the elementary teaching of Probability Calculus had been taught at the Polytechnic School of Lisbon, and in 1911 a quarterly chair of Probability Calculus was created, which in 1932 became an annual chair, taught by Victor Hugo de Lemos and Pedro Braumann (cf.
+Since 1865, a chair of Infinitesimal Calculus including the elementary teaching of Probability Calculus had been taught at the Polytechnic School of Lisbon, and in 1911 a quarterly chair of Probability Calculus was created at the Faculty of Sciences, University of Lisbon, which in 1932 became an annual chair, taught by Victor Hugo de Lemos and Pedro Braumann (cf.
 Fontes de Sousa, 2005 [51]).
-The lecture notes from Victor Hugo de Lemos’s course was edited in 1945 by J.
+The lecture notes [92] from Victor Hugo de Lemos’s course was edited in 1945 by J.
 Cabral Madeira.
+The five volumes of Adolpho Loureiro’s Os Portos Marítimos de Portugal e ilhas Adjacentes: Atlas [95] [The Sea Ports of Portugal and Adjacent Islands], containing many statistical tables and details, were published from 1904 to 1909.
   • [Uncaptioned image]
-  • As far as publications in the area of Probability are concerned, the first noteworthy event was the publication of a booklet intended to the layman, in 1904, volume 223 of the Biblioteca do Povo e das Escolas [Library for the People and the Schools], Noções sobre Cálculo das Probabilidades, Theoria dos Erros e Méthodo dos Mínimos Quadrados [Notions of Probability Calculus, Theory of Errors and the Least Squares Method], by Rodolpho Guimarães.
-As Rui Santos (2008a, 2024a, 2024b) [26, 27, 28] comments, Rodolpho Guimarães (9)(9)(9)  Rodolfo Guimarães (1866-1918) was an officer in the Portuguese army (Colonel) and a professor at the Army School of Lisbon (Military Academy).
+  • As far as publications in the area of Probability are concerned, the first noteworthy event was the publication of a booklet intended to the layman, in 1904, volume 223 of the Biblioteca do Povo e das Escolas [Library for the People and the Schools], Noções sobre Cálculo das Probabilidades, Theoria dos Erros e Méthodo dos Mínimos Quadrados [75] [Notions of Probability Calculus, Theory of Errors and the Least Squares Method], by Rodolpho Guimarães.
+As Rui Santos [150, 154, 155] comments, Rodolpho Guimarães888  Rodolfo Guimarães (1866-1918) was a Colonel in the Portuguese army and a professor at the Army School of Lisbon (Military Academy).
 He also served at the Astronomical Observatory of Tapada in Lisbon.
 Currently, Guimarães is best known for his research in the History of Mathematics, particularly for publishing a catalog of mathematical works by Portuguese authors during the 19th century and for promoting the work of Pedro Nunes (1502–1578), as well as some contributions in the field of Geometry.
-For detailed information on Rodolfo Guimarães, see Saraiva, L.
-(1997) [25].
+For detailed information on Rodolfo Guimarães, see Saraiva [157].
 , in the scant 64 pages of this small volume, manages to present an interesting introduction to the topic.
-The second notable publication is by José Freire de Sousa Pinto (1855-1911) — the Noções de Cálculo das Probabilidades para o Estabelecimento das Bases da Estatística [Notions of Probability Calculus for Establishing the Bases of Statistics], analysed in Santos (2008b, 2016) [29, 30].
+The second notable publication is by José Freire de Sousa Pinto (1855-1911) — the Noções de Cálculo das Probabilidades para o Estabelecimento das Bases da Estatística [173] [Notions of Probability Calculus for Establishing the Bases of Statistics], analysed in Santos [151, 153].
 José de Sousa Pinto, a professor at the University of Coimbra, discusses the use of Probability Calculus in modeling random phenomena.
-In this work, the author begins by explaining the process of constructing knowledge in any science, concluding that the advancement of knowledge results from the combination of observing the phenomenon and theoretical reasoning.
+In this work, the author begins by explaining the process of constructing knowledge in any experimental science, concluding that the advancement of knowledge results from the combination of observing the phenomenon and theoretical reasoning.
 Thus, any theory arises from the observation of a phenomenon, initially being an approximation.
@@
 It is in this successive chain of dual effort, observation and reasoning, that scientific knowledge advances towards a complete understanding of the phenomenon, where the formalized theory will always occur whenever the phenomenon happens.
-It should be noted that the analysis of prerequisites for teaching Statistics was a subject of reflections by Leite Pinto (1945), in Noções de Matemática necessárias ao estudo da Estatística [Notions of Mathematics Necessary for the Study of Statistics], and by Amaro Guerreiro, in Em Torno do Ensino da Estatística [About the Teaching of Statistics], 1952-53.
+It should be noted that the analysis of prerequisites for teaching Statistics was a subject of reflections by Leite Pinto [90], in Noções de Matemática necessárias ao estudo da Estatística [Notions of Mathematics Necessary for the Study of Statistics], and by Amaro Guerreiro, in Em Torno do Ensino da Estatística [73] [About the Teaching of Statistics], 1952-53.
   • [Uncaptioned image]
-  • In 1914, Diogo Pacheco de Amorim (1888-1976) defended his doctoral dissertation at the University of Coimbra on Elementos de Cálculo das Probabilidades [Elements of Probability Calculus], the aim of which was to establish rigorous foundations for the theory of Probability.
-In fact, this objective failed, and another doctoral thesis on Probability, by Manuel Reis (1929), is very critical of Pacheco de Amorim’s thesis.
+  • In 1914, Diogo Pacheco d’Amorim (1888-1976) defended his doctoral dissertation at the University of Coimbra on Elementos de Cálculo das Probabilidades [127] [Elements of Probability Calculus], the aim of which was to establish rigorous foundations for the theory of Probability.
+In fact, this objective failed, and another doctoral thesis on Probability [144], by Manuel Reis (1929), is very critical of Pacheco d’Amorim’s thesis.
   • [Uncaptioned image]
-Later, in a unique work by Luís de Albuquerque (1959) on the foundations of the Calculus of Probabilities, those dissertations are not even mentioned.
-Pacheco de Amorim’s thesis has been exhaustively studied by Rui Santos (2008c) [31].
-See, as a simple introduction, Pestana and Velosa (2001) [32] and Pestana, D. (1994) [33]. There is an annotated English translation, Mendonça et al. (2007) [34].
-In fact, it was only in 1933 that Kolmogorov (1933) [35] made an axiomatic construction of Probability, inspired by Maurice Fréchet’s ideas on the use of Measure Theory.
+Later, in a unique work by Luís de Albuquerque [2] on the foundations of the Calculus of Probabilities, those dissertations are not even mentioned.
+Pacheco d’Amorim’s thesis has been exhaustively studied by Rui Santos [152], see also [107].
+See, as a simple introduction to the work of Pacheco d’Amorim, Pestana and Velosa [139] and Pestana [138]. There is an annotated English translation of his doctoral thesis by Mendonça et al.  [107].
+In fact, it was only in 1933 that Kolmogorov [86] made an axiomatic construction of Probability, inspired by Maurice Fréchet’s ideas on the use of Measure Theory.
   • [Uncaptioned image]
-  • The Theory of Measurement was tackled by José Joaquim Dionísio (1957) (who, in what regards Probability, published a brief note on Shannon’s entropy in 1959 in Gazeta de Matemática), and became the topic of choice for Pedro Braumann (1969, 1978).
+  • Measure Theory was tackled by José Joaquim Dionísio [43] (who, in what regards Probability, published a brief note on Shannon’s entropy [44] in Gazeta de Matemática), and became the preferred field of work for Pedro Braumann [19, 20].
   • [Uncaptioned image]
-Pedro Braumann, who was a fellow at Stanford in 1955-56, was also the author of an excellent exposition on the central theme of classical probability theory, the arithmetic of probability laws, [Introduction to the Study of Limits of Sums of Independent Casual Variables] (1958).
+Pedro Braumann, who was a fellow at Stanford in 1955-56, was also the author of an excellent exposition on the central theme of classical Probability Theory, the arithmetic of probability laws, [Introduction to the Study of Limits of Sums of Independent Casual Variables] [18].
 Statistics at the University of Coimbra
-The teaching of Statistics descending from the links to Political Economy persisted at the Faculty of Law of the University of Coimbra, giving rise to the publication of various texts (Gouveia et al, 1913; Marnoco e Sousa(10)(10)(10)  The incomplete copy from FDUL did not allow us to access some parts of the book; it should be noted that it has a section on “the sceptics of Statistics”, a subject that is currently widely published.
+The teaching of Statistics associated to Political Economy persisted at the Faculty of Law of the University of Coimbra, giving rise to the publication of various texts (Gouveia et al.
+[67], Marnoco e Souza999  The incomplete copy from FDUL did not allow us to access some parts of the book; it should be noted that it has a section on “the sceptics of Statistics”, a subject that is currently widely discussed.
 In Chapter III (Statistical Laws), sections 59 and 60 are respectively “Law of large numbers” and “Law of small numbers”.
-Poisson, wanted his result on the convergence of a succession of binomials to Poisson to be known as the law of small numbers, renamed “law of large numbers” Jacques Bernoulli’s “golden theorem” on the convergence of arithmetic means towards the expectation.
-Only the expression “law of large numbers” persisted., 1913; Silva and Rocha, 1919; Rocha and Monteiro, 1923; Pereira, 1929).
+Poisson, wishing that his result on the convergence of a succession of binomials to Poisson to be known as the law of small numbers, renamed “law of large numbers” Jacques Bernoulli’s “golden theorem” on the convergence of arithmetic means towards the expectation.
+Only the expression “law of large numbers” persisted.
+[101], Silva and Rocha [162], Rocha and Monteiro [145], Pereira [135]).
 On the other hand, at the Anthropology Institute of the University of Coimbra, Eugénio Tamagnini had an innovative perspective on the contribution of Statistics in Biometrics research.
-In Revista da Faculdade de Ciências da Universidade de Coimbra, Tamagnini (1937) published an innovative work in Portugal on analysis of variance.
+In Revista da Faculdade de Ciências da Universidade de Coimbra, Tamagnini [181] published an innovative work in Portugal on analysis of variance.
 Tamagnini also hired W.
@@
 Seal”, and it can be conjectured that he had something to do with the British intelligence services during the war.
-Fonseca (2011) [38] states that “It is in this context of the necessity to improve the statistical treatment of anthropometric, physiological and demographic data, that one understands the arrival of W.
+Fonseca [50] states that “It is in this context of the necessity to improve the statistical treatment of anthropometric, physiological and demographic data, that one understands the arrival of W.
 L.
@@
 These are just some of the questions that (at least) for now, will remain without an answer.”.
-Between 1942 and 1945 Stevens published a remarkable set of works which he categorised as Questions of Method — namely on the $\chi^{2}$ significance test, discriminant analysis, statistical distributions, parameters estimation —, disseminating innovations in mathematical statistics by Karl Pearson, Ronald Fisher, Egon Pearson, Jerszy Neyman and their schools.
-Stevens also published, with Aureliano Quintanilha and Hugo Ribeiro (1942), an interesting article on the application of Probability Calculus in the investigation of non sexual reproduction by plasmogamy followed by karyogamy in the higher fungi basidiomycetes, in the Gazeta de Matemática (both in the index and in the article the initials W and L are interchanged).
+Between 1942 and 1945 Stevens published a remarkable set of works which he categorised as Questions of Method — namely on the $\chi^{2}$ significance test, discriminant analysis, statistical distributions, parameters estimation —, disseminating innovations in mathematical statistics by Karl Pearson, ‘Student’, Ronald Fisher, Egon Pearson, Jerszy Neyman and their schools.
+Stevens also published, with Aureliano Quintanilha and Hugo Ribeiro [141] an interesting article on the application of Probability Calculus in the investigation of non sexual reproduction by plasmogamy followed by karyogamy in the higher fungi basidiomycetes, in the Gazeta de Matemática (both in the index and in the article the initials W and L are interchanged).
 Statistics at the Technical University of Lisbon
@@
 Special mention should be made of
-  • Manuel Zaluar Nunes (1907-1967), who started as an assistant at ISCEF, before going to study in Paris as a fellow of the Instituto de Alta Cultura, published in 1933 with Mário Santos a concise exposition of Elementos de Cálculo das Probabilidades e de Estatística Matemática [Elements of Probability Calculus and Mathematical Statistics]: written in accordance with the program of the 2nd chair of the Instituto Superior de Ciências Económicas e Financeiras.
-Founder of the Gazeta de Matemática, Portugaliae Mathematica, and the Sociedade Portuguesa de Matemática [Portuguese Mathematical Society], he was an Assistant Professor at FCUL and later a Full Professor at ISA, a position from which he was dismissed in 1947 for political reasons, having to migrate first to France, where he was a research assistant at the Centre National de la Recherche Scientifique, and later to Brazil, where he was hired as a professor at the University of Recife.
-  • Francisco Paulo Pinto Leite (1902-2000), who was a lecturer at the Sorbonne from 1931 to 1933.
+  • Manuel Zaluar Nunes (1907-1967), who started as an assistant at ISCEF, before going to study in Paris as a fellow of the Instituto de Alta Cultura, published in 1933 with Mário Santos a concise exposition of Elementos de Cálculo das Probabilidades e de Estatística Matemática [149] [Elements of Probability Calculus and Mathematical Statistics]: written in accordance with the program of the 2nd chair of the Instituto Superior de Ciências Económicas e Financeiras.
+Founder of the Gazeta de Matemática, of Portugaliae Mathematica, and of the Sociedade Portuguesa de Matemática [Portuguese Mathematical Society], he was an Assistant Professor at FCUL and later a Full Professor at ISA, a position from which he was dismissed in 1947 for political reasons, having to migrate first to France, where he was a research assistant at the Centre National de la Recherche Scientifique, and later to Brazil, where he was hired as a professor at the University of Recife.
+  • Francisco Paulo Pinto Leite (1902-2000) was lecturer at Sorbonne from 1931 to 1933.
 One of the founders of the Portuguese Mathematical Society, he was appointed Full Professor at ISCEF in 1940, retiring in 1973.
-There are various versions of the course notes for the Statistics course he taught at ISCEF, and he was also the first lecturer, in 1952-53, of the chair of Econometrics, with Henri Guitton (University of Dijon) and José de Castañeda (University of Madrid), having as assistant Manuel Jacinto Nunes (further information in Machado and Santos Silva, 2002 [39]).
+There are various versions of the course notes for the Statistics course he taught at ISCEF, and he was also the first lecturer, in 1952-53, of the chair of Econometrics, with Henri Guitton (University of Dijon) and José de Castañeda (University of Madrid), having as assistant Manuel Jacinto Nunes (further information in Machado and Santos Silva [97]).
     It should be noted that 1953 saw the publication of Trabalhos do Seminário de Econometria Dirigido pelo Prof.
@@
 O.
-Wold and the translation of Wold’s book on dynamic systems by Jacinto Nunes, and that Mira Fernandes (who prefaced the book by Mário Santos and Zaluar Nunes) consistently supported research in Statistics and Econometrics, see the series of eight volumes of Estudos de Matemática, Estatística e Econometria [Studies in Mathematics, Statistics and Econometrics] that he edited between 1956 and 1964.
-  • José Sebastião e Silva (1914-1972), a fellow of the Instituto de Alta Cultura in Rome from 1942 to 1946, in 1951 won the public competition for the position of Full Professor at the Instituto Superior de Agronomia, taking charge of the chairs of General Mathematics and Calculus of Probabilities.
-For the latter chair, in 1954-55, he prepared didactic texts, collected in the course notes Cálculo das Probabilidades [Calculus of Probabilities] published at ISA, with an appendix from 1957-58 on correlation and regression, adjustment by the method of least squares, and Student’s $t$ and Fisher-Snedecor’s $F$ statistics, cf.
-Neves (2005) [40], pp.222–223.
-In 1991 Calouste Gulbenkian Foundation republished this work, now unfortunately out of print.
-Perhaps as a result of the influence of the work developed by Sir Ronald Fisher and his collaborators at Rothamsted, new areas of Statistics were developed at ISA, particularly in the area of Experimental Design (Costa, 1933; Carvalho, 1945, 1946) and Linear Models (Pestana, 1957).
-Professor Varennes e Mendonça taught Probabilities at ISA before the appointment of Professor Sebastião e Silva, publishing a brief course on Noções de Cálculo das Probabilidades [Notions of Probability Calculus] (1950), for details cf.
-Neves (2005) [40], pp.
+Wold [206] [Works presented at the Seminar of Econometry drected by Professor H.
+O.
+Wold] and the translation of Wold’s book on dynamic systems [206] by Jacinto Nunes, and that Mira Fernandes (who prefaced the book by Mário Santos and Zaluar Nunes) consistently supported research in Statistics and Econometrics, see the series of eight volumes of Estudos de Matemática, Estatística e Econometria [Studies in Mathematics, Statistics and Econometrics] [108] that he edited between 1956 and 1964.
+  • José Sebastião e Silva (1914-1972), a fellow of the Instituto de Alta Cultura in Rome from 1942 to 1946, in 1951 won the public competition for the position of Full Professor at the Instituto Superior de Agronomia, taking charge of the chairs of General Mathematics and of Calculus of Probabilities.
+In 1954-55, he prepared didactic texts, collected in the course notes Cálculo das Probabilidades[158] [Calculus of Probabilities] published at ISA, with an appendix from 1957-58 on correlation and regression, adjustment by the method of least squares, and Student’s $t$ and Fisher-Snedecor’s $F$ statistics, cf.
+Neves [125], pp.222–223.
+In 2000 Calouste Gulbenkian Foundation republished this work [159], now unfortunately out of print.
+Perhaps as a result of the influence of the work developed by Sir Ronald Fisher and his collaborators at Rothamsted, new areas of Statistics were developed at ISA, particularly in the area of Experimental Design (Costa [36], Carvalho [28, 29]) and Linear Models (Pestana [137]).
+  • Professor Varennes e Mendonça, who graduated in Agronomy Engineering at ISA, taught Probabilities at ISA before the appointment of Professor Sebastião e Silva, publishing a brief course on Noções de Cálculo das Probabilidades [194] [Notions of Probability Calculus], for details cf.
+Neves [125], pp.
 224–225.
-Varennes e Mendonça early on realized the difficulties of translating statistical terms, publishing an interesting article on terminology in 1942.
+Varennes e Mendonça early on realized the difficulties of translating statistical terms, publishing an interesting article [193] on terminology in 1942.
   • [Uncaptioned image]
-  • Murteira and Madureira (1958) supplied interesting translations, but it wasn’t until much later that the Specialised Committee on Statistical Nomenclature of the Portuguese Statistical Society and the Brazilian Statistical Association, whose members were Carlos Daniel Paulino (President); Dinis Pestana; João Branco; Lucília Carvalho; Alfredo Egídio dos Reis; Júlio Singer (Chairman); Lúcia Barroso; Wilton Bussab, prepared the Statistics Glossary English-Portuguese, Portuguese Statistical Society / Brazilian Statistical Association https://spestatistica.pt/glossario, which continues to evolve; see also Ventura, C.; Gomes, M.I.; Pestana, D.; Pestana, P.
-(2013).
-Glossary of Statistical Terms: German, French, English, Portuguese, INE, Lisbon.
+  • Murteira and Madureira [122] supplied interesting translations, but it wasn’t until much later that the Specialised Committee on Statistical Nomenclature of the Portuguese Statistical Society and the Brazilian Statistical Associationprepared the Statistics Glossary English-Portuguese [133], https://spestatistica.pt/glossario, which is periodically updated.
+See also Ventura [197] et al’s [Glossary of Statistical Terms: German, French, English, Portuguese].
 Publishers invest in Statistics — translations and dissemination
-The translation, in 1945, of Snedecor’s Statistical Methods Applied to Experiments in Agriculture and Biology certainly had a major impact on the development of Statistics in Portugal.
+The translation, in 1945, of Snedecor’s Statistical Methods Applied to Experiments in Agriculture and Biology [163] certainly had a major impact on the development of Statistics in Portugal.
 The popularisation of interest in Probability and Statistics, with an increase in the number of students taking courses in these areas, finally aroused the interest of publishers.
   • [Uncaptioned image]
-  • Publicações Europa-América pioneered this move to publish statistical books, with the publication of Gustavo de Castro’s books in the collection it Saber [Knowledge] (1959, 1965), an author chosen probably because he had previously published a book on Inferência Estatística [Statistical Inference] at LNEC [National Laboratory of Civil Engineering] in 1953.
+  • Publicações Europa-América pioneered this move to publish statistical books, with the publication of Gustavo de Castro’s books [32, 33] in the collection it Saber [Knowledge] (1959, 1965), an author chosen probably because he had previously published a book on Inferência Estatística [31] [Statistical Inference] at LNEC [National Laboratory of Civil Engineering] in 1953.
   • [Uncaptioned image]
-This was followed by the publication of the translation of Tippett’s Statistics (1955), and in 1973-75 Europa-América published the two volumes of P.
-Dagnelie’s Théorie et Méthodes Statistiques: Applications Agronomiques, translated by A.
+This was followed by the publication of the translation of Tippett’s Statistics [190] in 1955 (reprinted in 1968), and in 1973-75 Europa-América published the two volumes of P.
+Dagnelie’s Théorie et Méthodes Statistiques: Applications Agronomiques [37], translated by A.
 St.
@@
 L.
-Lévy’s Comprendre les Statistiques.
+Lévy’s Comprendre les Statistiques [96].
   • [Uncaptioned image]
   • [Uncaptioned image]
-  • Editorial Estúdios COR also ventured out with the publication in 1968 of a translation of Gnedenko and Khinchine’s book An Elementary Introduction to the Theory of Probability, with a preface and translation by António Simões Neto; and Despertar published a translation of one of the highest quality and most successful introductory books, Moroney’s Facts from Figures, in 1969.
+  • Editorial Estúdios COR also ventured out with the publication in 1968 of Gnedenko and Khinchine’s book Elementary Introduction to the Theory of Probability [58], with preface and translation by A.
+Simões Neto.
+Despertar published a translation of one of the highest quality and most successful introductory books, Moroney’s Facts from Figures [113], in 1969.
   • [Uncaptioned image]
 Some highlights
-The books by Guimarães (1904) and Sousa Pinto (1913) were analysed by Rui Santos, and those by Varennes e Mendonça (1950), Sebastião e Silva (1954) and Pestana (1957) by Manuela Neves [40], as already mentioned.
-Of the others, we think the publications by Santos and Zaluar Nunes (1933), Braumann (1958), Reis and Sarmento (1960) and Dias Agudo (1961) deserve special mention.
+The books by Guimarães [75] and Sousa Pinto [173] were analysed by Rui Santos, and those by Varennes e Mendonça [194], Sebastião e Silva [158] and Pestana [137] by Manuela Neves [125], as already mentioned.
+Of the others, we think the publications by Santos and Zaluar Nunes [149], Braumann [18], Reis and Sarmento [143] and Dias Agudo [42] deserve special mention.
 The publications by E.
-Tamagnini (1937) and W.
+Tamagnini [181] and W.
 L.
-Stevens (1942-45) in the Revista da Faculdade de Ciências da Universidade de Coimbra [Review of the Faculty of Sciences of Coimbra University] also deserve special mention.
-  • The book by Santos and Zaluar Nunes (1933) has a modernity that contrasts sharply with the 1932 book by Bueno y Martins, which has a perspective closer to official statistics.
+Stevens [174, 175, 176, 177, 178] in the Revista da Faculdade de Ciências da Universidade de Coimbra [Review of the Faculty of Sciences of Coimbra University] also deserve special mention.
+  • The 1933 book by Santos and Zaluar Nunes [149] has a modernity that contrasts sharply with the 1932 book by Bueno y Martins [23], which has a perspective closer to official statistics.
     • [Uncaptioned image]
-    • However, when it comes to descriptive statistics, Bueno y Martins’ book is more detailed, for example, in addition to the arithmetic mean, he considers geometric, harmonic and weighted averages; and in what regards graphical representations, Bueno y Martins discusses cartograms, stereograms, presenting various interesting pictograms, and discusses as well as association and contingency, and correlation.
+    • However, when it comes to descriptive statistics, Bueno y Martins’ book is more detailed, for example, in addition to the arithmetic mean, he considers geometric, harmonic and weighted averages.
     • [Uncaptioned image]
-    It should be noted, however, that some of the more elaborate graphics in Bueno y Martins (1932) come from other sources.
-For example, the extra text between pages 138 and 139 is a reproduction of the cover of the August 1930 issue of the Boletim da Direção Geral de Estatística [Bulletin of the General Directorate of Statistics], see the Catálogo Bibliográfico [Bibliographic Catalogue], p.19, published by INE in 2013.
+    In what regards graphical representations, Bueno y Martins discusses cartograms, stereograms, presenting various interesting pictograms, and discusses as well as association and contingency, and correlation.
+It should be noted, however, that some of the more elaborate graphics in Bueno y Martins [23] come from other sources.
+For example, the extra text between pages 138 and 139 is a reproduction of the cover of the August 1930 issue of the Boletim da Direção Geral de Estatística [Bulletin of the General Direction of Statistics], see the Catálogo Bibliográfico [Bibliographic Catalogue], p.19, published by INE in 2013 [82].
 Bueno y Martins’ main statistical reference is the monograph by Yule (the bibliographical sources used are indicated at the end of each section, which is very useful), and it is worth noting that Bueno y Martins indicates several books on Probability in the final bibliography, including A Treatise on Probability by Maynard Keynes.
@@
 Amaral Turkman, M. A.; Silva, G. (2000) Modelos Lineares Generalizados — da Teoria à Prática [Generalized Linear Models — from Theory to Practice]
-Brilhante, M. F.; Pestana, D.; Rocha, J.; Rocha, M. L.; Velosa, S. (2001) 2ª ed. 2011, Inferência Estatística sobre a Localização Usando a Escala [Statistical Inference about Location Using Scale]
+Brilhante, M. F.; Pestana, D.; Rocha, J.; Rocha, M. L.; Velosa, S. (2001) 2ª ed. 2011, Inferência Estatística sobre a Localização Usando a Escala [Statistical Inference about Location Using the Scale]
 Muller, D. (2002) Modelos Heterocedásticos. Aplicações com o software Eviews [Heteroscedastic Models. Applications with Eviews software]
@@
 Amaral Turkman, M. A.; Paulino, C. D. (2015) Estatística Bayesiana Computacional — uma Introdução [Computational Bayesian Statistics — an Introduction]
-Brilhante, M. F. (2017) An Introduction to Meta-Analysis [An Introduction to Meta-Analysis]
-Amado, C.; Nunes, C.; Sardinha, A. (2019) Análise Estatística de Dados Financeiros [Statistical Analysis of Financial Data]
+Brilhante, M. F. (2017) Uma Introdução à Meta-Análise [An Introduction to Meta-Analysis]
+Amado, C.; Nunes, C.; Sardinha, A. (2019) Análise Estatística de Dados Financeiros [Statistical Analysis of Finance Data]
 Afonso, P. M. (2023). Modelação Conjunta de Dados Longitudinais e de Sobrevivência [Joint Modeling of Longitudinal and Survival Data]
 Before the SPE Congresses, CEAUL organized conferences, whose proceedings were published:
-Actas I Colóquio de Estatística e Investigação Operacional, Lisboa, Centro de Estatística e Aplicações, 1977.
-Actas II Colóquio Estatística e Investigação Operacional, Fundão/Covilhã, CEAUL/SPEIO, 1981.
-Actas III Colóquio Estatística e Investigação Operacional, CEAUL/SPEIO, Tróia, 1985.
+Actas, I Colóquio de Estatística e Investigação Operacional, Lisboa, Centro de Estatística e Aplicações, 1977.
+Actas, II Colóquio Estatística e Investigação Operacional, Fundão/Covilhã, CEAUL/ SPEIO, 1981.
+Actas, III Colóquio Estatística e Investigação Operacional, CEAUL/SPEIO, Tróia, 1985.
 Proceedings of the SPE Congresses:
@@
 Manuela Souto de Miranda e Isabel Pereira.
-Afirmar a Estatística: um Desafio para o Século XXI [Affirming Statistics: a Challenge for the 21st Century], Atas do VI Congresso da SPE, 1999, Tomar; ed.
+Afirmar a Estatística: um Desafio para o Século XXI [Asserting Statistics: a Challenge for the 21st Century], Atas do VI Congresso da SPE, 1999, Tomar; ed.
 Carlos Daniel Paulino, António Pacheco, Ana Pires e Ferreira da Cunha.
-Um Olhar sobre a Estatística [A Look at Statistics], Actas do VII Congresso da SPE, 2001, Ofir; ed.  Pedro Oliveira e Emília Athayde.
+Um Olhar sobre a Estatística [A Look upon Statistics], Actas do VII Congresso da SPE, 2001, Ofir; ed.  Pedro Oliveira e Emília Athayde.
 A Estatística em Movimento [Statistics in Motion], Actas do VIII Congresso da SPE, 2001, Peniche; ed.  Maria Manuela Neves, Jorge Cadima, Maria João Martins e Fernando Rosado.
@@
 Estatística com Acaso e Necessidade [Statistics with Chance and Necessity], Actas do XI Congresso da SPE, 2004, Faro; ed.  Paulo M. M. Rodrigues, Efigénio da Luz Rebelo e Fernando Rosado.
-Estatística Jubilar [Jubilee Statistics], Actas do XII Congresso Anual da SPE, 2005, Évora; ed.  Carlos Braumann, Paulo Infante, Manuela M. Oliveira, Russell Alpízar-Jara e Fernando Rosado.
+Estatística Jubilar [Jubilar Statistics], Actas do XII Congresso Anual da SPE, 2005, Évora; ed.  Carlos Braumann, Paulo Infante, Manuela M. Oliveira, Russell Alpízar-Jara e Fernando Rosado.
 A Ciência Estatística [Statistical Science], Actas do XIII Congresso da SPE, 2006, Ericeira; ed.
@@
 Fernanda Oliveira, Margarida Mendes Leal e Fernando Rosado.
-Estatística, Ciência Interdisciplinar [Statistics, Interdisciplinary Science], Actas do XIV Congresso da SPE, 2007, Covilhã; ed.  Maria Eugénia Ferrão, Célia Nunes e Carlos Braumann.
+Estatística, Ciência Interdisciplinar [Statistics, an Interdisciplinary Science], Actas do XIV Congresso da SPE, 2007, Covilhã; ed.  Maria Eugénia Ferrão, Célia Nunes e Carlos Braumann.
 Estatística: da Teoria à Prática [Statistics: from Theory to Practice], Actas do XV Congresso da SPE, 2008, Lisboa; ed.
@@
 Dias, Maria de Fátima Salgueiro, Helena Carvalho, Paula Vicente e Carlos Braumann.
-Estatística: Arte de Explicar o Acaso [Statistics: Art of Explaining Chance], Actas do XVI Congresso Anual da SPE, 2009, Vila Real; ed.
+Estatística: Arte de Explicar o Acaso [Statistics: the Art of Explaining Chance], Actas do XVI Congresso Anual da SPE, 2009, Vila Real; ed.
 Irene Oliveira, Elisabete Correia, Fátima Ferreira, Sandra Dias e Carlos Braumann.
@@
 New Frontiers in Statistics and Data Science, Springer Proceedings in Mathematics & Statistics, Springer, XXVI Congresso da SPE, 2024, Guimarães (ed.
-Lígia Henriques-Rodrigues, Raquel Menezes, Luís Meira Machado, Miguel de Carvalho e Susana Faria), in print.
+Lígia Henriques-Rodrigues, Raquel Menezes, Luís Meira Machado, Miguel de Carvalho e Susana Faria).
 Appendix 2
-Index of the Manual de Estatística Médica (Manual of Medical Statistics) by Carlos Santos Reis and Alexandre Sarmento.
+Table of Contents of the Manual de Estatística Médica [143] [Manual of Medical Statistics] by Carlos Santos Reis and Alexandre Sarmento.
 Introduction 3
@@
 1 – The collection or collection of data 32
-      1.1 – Harvest method 33
+        1.1 – Collection method 33
         1.2 – Observation methods 34
@@
         1.4 – Investigation modalities 37
-      1.5 – Source of data 42
+        1.5 – Sources of data 42
         1.6 – Preparation of questionnaires 43
@@
         1.8 – Sources of errors in surveys 48
-      1.9 – Observation record 52
+        1.9 – Observations record 52
 6 – Data verification 56
@@
 2 – Quantitative characters 126
-11 – Position measurements 129
+11 – Location measurements 129
 1 – Median 130
 2 – Arithmetic mean 133
-   3 – Fashion 137
+3 – Mode 137
 4 – Geometric mean 138
-   5 – Comparison of main position measurements 142
+5 – Comparison of the main location measurements 142
 12 – Dispersion and shape measurements 145
@@
 13 Simplified calculation of parameters 154
-   1 – Fashion 155
+1 – Mode 155
 2 – Median 156
@@
 1 – Estimation and confidence limits of the mean 207
-   2 – Estimation and confidence limits of variance 210
+2 – Estimation and confidence limits of the variance 210
 3 – Estimation and confidence limits of a proportion 210
@@
 1 – Conformity tests for parameters 219
-   2 – Distribution Conformance Tests 220
+2 – Distribution of conformance tests 220
 19 – Homogeneity problem 232
@@
 16 – Correlation of the number of cases 376
-   17 – Position correlation 377
+17 – Location correlation 377
 18 – Conclusions 378
@@
 3 – Linear adjustment 388
-   4 – Curvy Fit 393
+4 – Curve fitting 393
 5 – Adjustment to theoretical distributions 400
@@
 […]
-Method    1 – Principles of the 420 method
+1 – Principles of the method 420
 2 – Checking the normality of a distribution 424
@@
 2 – Movements of a chronological series 434
-   3 – Guidance to follow when studying these series 434
+3 – Guidance for studying series 434
 4 – Data preparation 435
@@
 Appendix 3
-Index of Introdução à Estatística (Introduction to Statistics), notes for the course on General Mathematics for the degrees in Biological Sciences, Geological Sciences, and Adjunct Professors, based on the lectures of Professor Dias Agudo, compiled by Moreira Campos / Waldemar Nunes, Lisbon 1961/62
+Table of Contents of Introdução à Estatística [42] [Introduction to Statistics], notes for the course on General Mathematics for the degrees in Biological Sciences, Geological Sciences, and Adjunct Professors, based on the lectures of Professor Dias Agudo, compiled by Moreira Campos / Waldemar Nunes, Lisbon 1961/62
 Chapter I – Fundamental Concepts of Probability Theory
@@
                Incompatible and independent events
-      1.6 – Problem with repeated tests.
+        1.6 – Problem with repeated trials.
                Binomial formula
@@
                Binomial Distribution
-Chapter IV – Continuous Distributions*
+Chapter IV – Continuous Distributions
         4.1 – Probability density
@@
         8.4 – Comparison of two populations
+Acknowledgments
+This work is partially financed by national funds through FCT — Fundação para a Ciência e a Tecnologia under the project UIDB/00006/2020 https://doi.org/10.54499/UIDB/00006/2020.
+The authors are thankful to Prof.
+Antónia Amaral Turkman, Maria de Fátima Brilhante and Sandra Mendonça for useful observations on the original draft.
+The authors also extend their gratitude to the reviewers for their stimulating comments, corrections and informations that helped to improve the paper.